### v1

# Triality and Adjoint lifting for ${\rm GL}_{3}$

Triality and Adjoint lifting for ${\rm GL}_{3}$
Wee Teck Gan Department of Mathematics, National University of Singapore, 10 Lower Kent Ridge Road Singapore 119076 matgwt@nus.edu.sg
2010 Mathematics Subject Classification:
Primary 11F70, Secondary 22E50
1. Introduction
This paper and its prequel [ChG] aim to highlight some implications of the principle of triality in the Langlands program. The principle of triality refers to the fact that the (split) simply-connected ${\rm Spin}_{8}$ and the adjoint ${\rm PGSO}_{8}$ possess an order 3 outer automorphism $\theta$. In [ChG], by composing known cases of Langlands functorial lifting with the triality automorphism, we showed that one can obtain new and dramatically different instances of functorial lifting. In the present paper, we shall consider the implication of triality in the theory of twisted endoscopy.
After recalling some necessary background on the triality automorphism in §2, we determine the twisted endoscopic groups associated to the twisted space $({\rm PGSO}_{8},\theta)$ in §3. It turns out that the twisted endoscopic groups of $({\rm PGSO}_{8},\theta)$ are one of the following: the exceptional group ${\rm G}_{2}$, ${\rm SL}_{3}$ and ${\rm SO}_{4}$. For the case of $H={\rm SL}_{3}$, for example, the associated map of dual groups is the unique lift of the adjoint representation
$\tilde{{\rm Ad}}:H^{\vee}={\rm PGL}_{3}({\mathbb{C}})\longrightarrow{\rm Spin}_{8}({\mathbb{C}}).$
Indeed, $\tilde{\rm Ad}$ fits into the folllowing composite of morphisms of dual groups:
(1.1) $\begin{CD}{\rm GL}_{3}({\mathbb{C}})@>{}>{}>{\rm PGL}_{3}({\mathbb{C}})@>{\tilde{{\rm Ad}}}>{}>{\rm Spin}_{8}({\mathbb{C}})@>{}>{}>{\rm SO}_{8}({\mathbb{C}})@>{\rm std}>{}>{\rm GL}_{8}({\mathbb{C}}),\end{CD}$
which is nothing but the adjoint representation of ${\rm GL}_{3}$ on its semisimple Lie algebra $\mathfrak{sl}_{3}$.
Using the stable twisted trace formula established by Moeglin and Waldspurger, one would thus expect to show the (weak) twisted endoscopic lifting from ${\rm SL}_{3}$ to ${\rm PGSO}_{8}$. On the other hand, the functorial lifting associated to the first and third arrows are simply the pullbacks of automorphic forms from ${\rm GL}_{3}$ to ${\rm SL}_{3}$ and from ${\rm PGSO}_{8}$ to ${\rm SO}_{8}$ respectively, while the functorial lifting associated to the standard representation of ${\rm SO}_{8}$ is known by [A, CKPSS]. Hence, by combining these known functorial liftings with the twisted endoscopic lifting induced by $\tilde{{\rm Ad}}$, one would be able to show the (weak) adjoint lifting from ${\rm GL}_{3}$ to ${\rm GL}_{8}$. Without analyzing the stable trace formula too deeply, we carry out this strategy in this paper and show the anticipated adjoint lifting under a mild local condition:
Theorem 1.2.
Assume that $\pi$ is a cuspidal representation of ${\rm GL}_{3}$ over a number field, such that for some finite place $v_{0}$, the local component $\pi_{v_{0}}$ is a discrete series representation. Then the weak adjoint lifting ${\rm Ad}(\pi)$ exists as an automorphic representation of ${\rm GL}_{8}$.
We recall that the adjoint lifting for ${\rm GL}_{2}$ is a classical result of Gelbart and Jacquet [GJ]. The local condition in the theorem can no doubt be removed by a more extensive investigation of the stable twisted trace formula, but we do not pursue this here, as it is necessarily much more technical. The general case will be handled in ongoing joint work with S.W. Shin and Y. Takanashi, the goal of which is to obtain an endoscopic classification for the group $G_{2}$ via the twisted endoscopy for $({\rm PGSO}_{8},\theta)$.
As a second result of this paper, we analyze the possible shapes of the isobaric decomposition of the automorphic representation ${\rm Ad}(\pi)$ supplied by Theorem 1.2; this is often referred to as a cuspidality criterion for the adjoint lifting. The results, contained in Theorems 8.2, 8.4, 9.2 and 10.1, can be summarized as follows:
Theorem 1.3.
Let $\pi$ be a unitary cuspidal automorphic representation of ${\rm GL}_{3}$ and assume that ${\rm Ad}(\pi)$ has been constructed as the composite of the Langlands functorial lifting given in (1.1), as discussed above. Then ${\rm Ad}(\pi)$ is a self-dual automorphic representation of orthogonal type with trivial central character. Moreover, one has:
(i) ${\rm Ad}(\pi)$ contains a cuspidal ${\rm GL}_{1}$-summand if and only if $\pi$ is obtained by automorphic induction from a cyclic cubic field extension. In this case, then isobaric deocmposition of ${\rm Ad}(\pi)$ is of type $1+1+3+3$ or $1^{8}$, depending on whether $\pi$ can be obtained as an automorphic induction from one or four cyclic cubic extensions.
(ii) ${\rm Ad}(\pi)$ contains a cuspidal ${\rm GL}_{2}$-summand if and only if $\pi$ is obtained by automoprhic induction from a non-cyclic cubic field extension. In this case, the isobaric decomposition of ${\rm Ad}(\pi)$ is of type $2+6$, $2+3+3$ or $2^{4}$, depending on whether $\pi$ can be obtained as an automorphic induction from one (for the first two cases) or four (for the last case) non-cyclic cubic extensions.
(iii) Assume that $\pi$ is not an automorphic induction from any cubic field extension, so that ${\rm Ad}(\pi)$ does not contain any cuspidal ${\rm GL}_{1}$- or ${\rm GL}_{2}$-summands. Then ${\rm Ad}(\pi)$ contains a cuspidal ${\rm GL}_{3}$-summand if and only if $\pi$ is a twist (by an automorphic character) of a self-dual cuspidal representation. In this case, the isobaric decomposition of ${\rm Ad}(\pi)$ is of type $3+5$.
(iv) Assume that we are not in the setting of (i), (ii) or (iii). Then one of the following holds:
  • (a)
    ${\rm Ad}(\pi)$ is the sum of two self-dual cuspidal ${\rm GL}_{4}$-summands of orthogonal type and nontrivial central character;
  • (b)
    ${\rm Ad}(\pi)$ is cuspidal.
Except in (i), ${\rm Ad}(\pi)$ is a discrete orthogonal L-parameter, i.e. it is the functorial lift of a globally generic cuspidal automorphic representation of the split group ${\rm SO}_{8}$. Moreover,
${\rm ord}_{s=1}L^{S}(s,{\rm Ad}(\pi),\wedge^{3})={\rm ord}_{s=1}L^{S}(s,{\rm Ad}(\pi),{\rm Sym}^{2})<0,$
so that the order of poles at $s=1$ of $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ (which is an L-function of Langlands-Shahidi type) is equal to the number of summands in the isobaric decomposition of ${\rm Ad}(\pi)$.
The proof of this theorem relies crucially on the functorial lifting results of Gelbart-Jacquet [GJ], Arthur-Clozel [AC], Ramakrishnan [Ra1], Kim-Shahidi [KS1] and Kim [Ki], as well as the cuspidality criteria for some of these liftings due to Ramakrishnan-Wang [RW] and Asgari-Raghuram [AR].
It is of course of great interest to determine precisely the fibers and image of the adjoint lifting. In particular, the question of the fibers of the adjoint lifting is closely related to the question of multiplicities and rigidity for the cuspidal spectrum of ${\rm SL}_{3}$. It has been shown by Blasius [B] that the multiplicity-one property and the rigidity of L-packets both fail for ${\rm SL}_{n}$ for $n\geq 3$. In other words, there are cuspidal representations of ${\rm SL}_{n}$ ($n\geq 3$) with multiplicity >1>1, and there are inequivalent cuspidal L-packets which are nearly equivalent. Lapid [L1] extended this line of study with a more quantitative analysis, especially for the endoscopic L-packets of ${\rm SL}_{n}$ obtained by endoscopic transfer (== automorphic induction) from elliptic tori. In particular, he showed that in the context of ${\rm SL}_{3}$, the so-called global multiplicity of a cuspidal endoscopic L-packet is either $1$ or $2$ (see [L1, §2, Pg 162] for the definition and [L1, Theorem 4] for the result). For example, in the context of Theorem 1.3(i), the degenerate case where ${\rm Ad}(\pi)$ is the sum of 8 cubic characters coincides precisely with the case where the global multiplicity is $2$.
Acknowledgments: We thank Gaetan Chenevier (who should have been a coauthor of this paper), Erez Lapid, Dipendra Prasad, Sug Woo Shin and Yugo Takanashi for helpful discussions at various stages of this work and comments on earlier drafts of this paper. In addition, we thank Henry Kim for sending us a copy of the paper [KS3] and Freydoon Shahidi for discussion on the fibers of the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ to ${\rm GL}_{6}$. The author is supported by a Tan Chin Tuan Centennial Professorship at NUS.
2. Triality
In this section, we recall some necessary background on the triality automorphism of ${\rm Spin}_{8}$ and ${\rm PGSO}_{8}$.
2.1. Symmetric composition algebras and ${\rm Spin}_{8}$
Let us first recall from [ChG, §2] a construction of ${\rm Spin}_{8}$ with its triality automorphism via the theory of symmetric composition algebras. For the definition and properties of symmetric composition algebras, the reader may consult [KMRT]. By [KMRT, Theorem 34.37], there are essentially two isomorphism classes of split symmetric composition algebra over a field $F$. These may be described as follows:
  • Start with the split octonion algebra $(\mathbb{O},\cdot,N)$ over $F$, where $N$ is its norm form permitting composition: $N(x\cdot y)=N(x)\cdot N(y)$ for all $x,y\in\mathbb{O}$. Then one defines a new multiplication on $\mathbb{O}$ by $x\ast y=\bar{x}\cdot\bar{y}$. The algebra $(\mathbb{O},\ast,N)$ is the para-octonion algebra and one still has $N(x\ast y)=N(x)\cdot N(y)$ for all $x,y\in\mathbb{O}$. Moreover, ${\rm Aut}(\mathbb{O},\cdot)={\rm Aut}(\mathbb{O},\ast)\simeq G_{2}^{\mathbb{O}}$.
  • Assume now that $F$ contains a primitive 3rd root of unity $\omega$ (so ${\rm char}(F)\neq 3$). The other split symmetric composition algebra of dimension $8$ is realised on the space $\mathbb{M}$ of trace zero $3\times 3$ matrices with multiplication given (miraculously) by [KMRT, Pg. 472, (34.18)]
    $x\ast y=\omega xy+(1-\omega)yx-\frac{1}{3}T(yx),\quad\text{ for $x,y\in\mathbb{M}$,}$
    where $T$ is the trace of a $3\times 3$ matrix. If the characteristic polynomial of $x\in\mathbb{M}$ is
    $P_{x}(t)=t^{3}+S(x)t-\det(x),$
    then the quadratic form on $\mathbb{M}$ which permits composition is $\frac{1}{3}\cdot S(x)$ [KMRT, Prop. 34.19]. The automorphism group of $(\mathbb{M},\ast),S$ is the group ${\rm PGL}_{3}\subset{\rm SO}(\mathbb{M},S)$ acting on $\mathbb{M}$ by conjugation. In other words, the action of ${\rm PGL}_{3}$ on $\mathbb{M}$ is isomorphic to the adjoint representation.
Thus, over a general field $F$ with ${\rm char}(F)\neq 3$, the $F$-forms of the above symmetric composition algebras are classified by $H^{1}(F,G_{2})$ and $H^{1}(F,{\rm PGL}_{3})$ respectively.
Given a symmetric composition algebra $(V,\ast,Q)$ of dimension $8$ over $F$, let us define
${\rm Spin}(V,\ast,Q)=\{(g_{1},g_{2},g_{3})\in{\rm SO}(V,Q)^{3}:g_{1}(x\ast y)=g_{2}(x)\ast g_{3}(y)\text{ for $x,y\in V$}\}.$
Then this group is known to be isomorphic to ${\rm Spin}_{8}$ over $\overline{F}$. Moreover, the cyclic permutation of the three coordinates in ${\rm SO}(V,Q)^{3}$ preserves the subgroup ${\rm Spin}(V,\ast,Q)$ and defines an order $3$ automorphism
$\theta_{V}:{\rm Spin}(V,\ast,Q)\longrightarrow{\rm Spin}(V,\ast,Q).$
The fixed group of $\theta_{V}$ is thus the subgroup ${\rm Aut}(V,\ast,Q)\subset{\rm SO}(V,Q)$. Because the automorphism $\theta$ necessarily preserves the center $Z_{{\rm Spin}(V,\ast,Q)}$, it descends to an automorphism of the adjoint group ${\rm PGSO}(V,Q)$, still denoted by $\theta_{V}$.
2.2. Splitting of $\rm Aut$}-sequence.
Recall that one has a short exact sequence
(2.1) $\begin{CD}1@>{}>{}>{\rm PGSO}_{8}@>{}>{}>{\rm Aut}({\rm Spin}_{8})@>{}>{}>{\rm Out}({\rm Spin}_{8})=S_{3}@>{}>{}>1.\end{CD}$
The data of a symmetric composition algebra $(V,\ast,Q)$ thus gives a splitting of this short exact sequence over the order 3 subgroup $A_{3}\subset S_{3}$. Since we could apply the above discussion to the two symmetric composition algebra $(\mathbb{O},\ast,N)$ and $(\mathbb{M},\ast,S)$ (when $F=\overline{F}$ say), we obtain two splittings of this short exact sequence over $A_{3}$ whose fixed subgroups are ${\rm G}_{2}^{\mathbb{O}}$ and ${\rm PGL}_{3}$ respectively. According to [KMRT, Pg. 487, Cor. 35.10 and Prop. 36.14], these are the only two possible splittings (up to inner conjugation) over $\overline{F}$.
In particular, the embedding
${\rm PGL}_{3}\simeq{\rm Aut}(\mathbb{M},\ast,S)\hookrightarrow{\rm Spin}(\mathbb{M},\ast,S)$
is invariant under the action of $\theta_{\mathbb{M}}$ and is the map denoted by $\tilde{\rm Ad}$ in the introduction.
2.3. Representations
By construction, the group ${\rm Spin}(V,\ast,Q)$ is equipped with three projections
$\rho_{j}:{\rm Spin}(V,\ast,Q)\longrightarrow{\rm SO}(V,Q),$
with ${\rm Ker}(\rho_{j})\simeq\mu_{2}\subset Z_{{\rm Spin}(V,\ast,Q)}$. Thus one has 3 inequivalent $8$-dimensional irreducible orthogonal representations of ${\rm Spin}(V,\ast,Q)$ which are permuted by $\theta$. Likewise, the projection from the three ${\rm SO}(V,Q)$’s gives 3 maps
$f_{j}:{\rm SO}(V,Q)\longrightarrow{\rm PGSO}(V,Q),$
which are permuted by $\theta$.
The following lemma will only be used much later in the paper, but this seems a convenient place to mention it:
Lemma 2.2.
. Let
$\rho_{1},\,\rho_{2},\,\rho_{3}:{\rm Spin}_{8}({\mathbb{C}})\longrightarrow{\rm SO}_{8}({\mathbb{C}})\subset{\rm GL}_{8}({\mathbb{C}})$
be the three inequivalent 8-dimensional irreducible representations of ${\rm Spin}_{8}({\mathbb{C}})$. Then
$\rho_{1}\otimes\rho_{2}\simeq\rho_{3}\oplus\wedge^{3}\rho_{3}.$
Now consider, for each $j$, the composite
$\begin{CD}{\rm GL}_{3}({\mathbb{C}})@>{}>{}>{\rm PGL}_{3}({\mathbb{C}})@>{\tilde{\rm Ad}}>{}>{\rm Spin}_{8}({\mathbb{C}})@>{\rho_{j}}>{}>{\rm SO}_{8}({\mathbb{C}})\subset{\rm GL}_{8}({\mathbb{C}}).\end{CD}$
Since $\tilde{\rm Ad}$ is invariant under the action of $\theta_{\mathbb{M}}$, this composite is independent of $j$ as a representation of ${\rm GL}_{3}({\mathbb{C}})$. As mentioned in the introduction, it is simply the adjoint representation $\rm Ad$} of ${\rm GL}_{3}({\mathbb{C}})$. The above lemma thus implies the following corollary:
Corollary 2.3.
One has an isomorphism
${\rm Ad}\otimes{\rm Ad}\simeq{\rm Ad}\oplus\wedge^{3}{\rm Ad}.$
of representations of ${\rm GL}_{3}({\mathbb{C}})$.
2.4. Chevalley-Steinberg épinglage.
For the purpose of this paper, the above geometric algebra approach to $({\rm PGSO}_{8},\theta)$ is less useful. Instead, we will give a root theoretic description of triality, by fixing a Chevalley-Steinberg system of épinglage for ${\rm PGSO}_{8}$. More precisely, we start with a Chevalley model of the split group ${\rm PGSO}_{8}$ over $\mathbb{Z}$ and fix a pair $T\subset B$ consisting of a maximal split torus contained in a Borel subgroup over $\mathbb{Z}$}. This gives a system of simple roots
$\Delta(T,B)=\{\alpha_{0},\alpha_{1},\alpha_{2},\alpha_{3}\},$
with $\alpha_{0}$ corresponding to the branch vertex in the Dynkin diagram:
${\alpha_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\alpha_{0}}$${\alpha_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\alpha_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
We may fix a pinning $(T,B,\{x_{\alpha_{i}}\})$, with $x_{\alpha_{i}}$ a basis vector in the root subspace $\mathfrak{u}_{\alpha_{i}}$. This then defines a splitting of the exact sequence
$\begin{CD}1@>{}>{}>{\rm PGSO}_{8}@>{}>{}>{\rm Aut}({\rm PGSO}_{8})@>{}>{}>{\rm Out}({\rm PGSO}_{8})@>{}>{}>1,\end{CD}$
with image
${\rm Aut}({\rm PGSO}_{8},T,B,\{x_{\alpha_{i}}\})\subset{\rm Aut}({\rm PGSO}_{8}).$
Since ${\rm Out}({\rm PGSO}_{8})\simeq S_{3}$, this gives us a triality automorphism $\theta$ which permutes $\{x_{\alpha_{1}},x_{\alpha_{2}},x_{\alpha_{3}}\}$ cyclically and fixes $x_{\alpha_{0}}$, so that
${\rm PGSO}_{8}^{\theta}\simeq{\rm G}_{2}.$
The dual group ${\rm Spin}_{8}({\mathbb{C}})$ thus inherits $T^{\vee}\subset B^{\vee}$ and a pinning $\{x_{\alpha_{i}^{\vee}}\}$ permuted by $\theta^{\vee}$ as well. Moreover,
${\rm Spin}_{8}({\mathbb{C}})^{\theta^{\vee}}\simeq{\rm G}_{2}({\mathbb{C}})$
with maximal torus $(T^{\vee})^{\theta^{\vee}}$. The action of $(T^{\vee})^{\theta^{\vee}}$ on $\mathfrak{spin}_{8}={\rm Lie}({\rm Spin}_{8}({\mathbb{C}}))$ gives rise to a ${\rm G}_{2}$ root system $\Phi_{G_{2}}$, which is simply the restriction of the roots of ${\rm Spin}_{8}({\mathbb{C}})$ (== coroots of ${\rm PGSO}_{8}$) to $(T^{\vee})^{\theta}$. Writing
$\alpha=\text{restriction of $\alpha_{i}^{\vee}$ ($i=1,2,3$)}\quad\text{and}\quad\beta=\text{restriction of $\alpha_{0}^{\vee}$},$
one has a system of simple roots $\{\alpha,\beta\}$ for $\Phi_{G_{2}}$. The long root spaces have dimension $1$ whereas the short root spaces have dimension $3$. Moreover, the triality automorphism $\theta^{\vee}$ preserves each of these root spaces for $\Phi_{G_{2}}$.
This Chevalley-Steinberg system of pinning provides us with a coordinate system with respect to which we will perform some explicit computations in the next section.
3. Twisted Endoscopic Groups
We shall be considering the theory of twisted endoscopy developed by Langlands and Kottwitz-Shelstad [KSh] to the twisted space defined by the triality automorphism $\theta$ of ${\rm PGSO}_{8}$. The general theory of the twisted trace formula was developed in [LW] and its stabilization has been established in the 2-volume book [MW1, MW2] of Moeglin-Waldspurger (modulo the twisted weighted fundamental lemma for general quasi-split groups). To exploit this stable twisted trace formula, we need to work out the elliptic twisted endoscopic data for $({\rm PGSO}_{8},\theta)$. This is the purpose of this section.
3.1. Lie algebra computations.
Our aim now is to determine, up to isomorphism, the elliptic endoscopic data $(H,s,\xi)$ of the twisted space $({\rm PGSO}_{8},\theta)$ over the number field $k$, following [KSh, §2.1]. [^1] In our setting, the cocycle $\bf a$} loc. cit. is chosen to be $1$, which forces the $a^{\prime}$ in (2.1.4a) to be $1$ as well. As ${\rm PGSO}_{8}$ is adjoint and split over $k$, their description simplifies a little and we may assume that these triple $(H,s,\xi)$ are as follows:
  • the element $s$ is a finite order element of $T^{\vee}$,
  • the (endoscopic) group $H$ is a split connected reductive group over $k$ equipped with an isomorphism between $H^{\vee}$ and the centralizer ${\rm C}_{s}$ in ${\rm Spin}_{8}({\mathbb{C}})$ of the element $s\cdot\theta^{\vee}\in{\rm Spin}_{8}({\mathbb{C}})\cdot\theta^{\vee}$,
  • (elliptic condition) $Z(\xi(H^{\vee}))^{0}\subset Z_{{\rm Spin}_{8}}$, or equivalently, $H$ is semisimple.
Indeed, it follows from a classical result of Steinberg [KSh, Chap. 1, §1.1, Pg. 14] that ${\rm C}_{s}$ is connected for all $s\in T^{\vee}$. It is harmless to view $\xi$ as an inclusion and write $H^{\vee}={\rm C}_{s}$.
To determine the possible $H^{\vee}$, we take a general $s\cdot\theta^{\vee}$, consider its adjoint action on $\mathfrak{spin}_{8}$ and determine the fixed Lie subalgebra $\mathfrak{spin}_{8}^{s\cdot\theta^{\vee}}$. Since we are interested in $s\cdot\theta^{\vee}$ up to conjugacy by ${\rm Spin}_{8}({\mathbb{C}})$, we may consider $s$ up to $\theta$-conjugacy by elements of $T^{\vee}({\mathbb{C}})$. As the group $T^{\vee}({\mathbb{C}})$ is divisible, it is not hard to see that one can assume that
$s\in T^{\vee}({\mathbb{C}})^{\theta^{\vee}}.$
Since ${\rm Spin}_{8}({\mathbb{C}})$ is simply-connected, $X_{*}(T^{\vee})=X^{*}(T)$ is equal to the root lattice of ${\rm PGSO}_{8}$. Hence we may write any element of $T^{\vee}({\mathbb{C}})^{\theta^{\vee}}$ as:
(3.1) $s(u,t)=\alpha_{0}(u)\cdot\alpha_{1}(t)\cdot\alpha_{2}(t)\cdot\alpha_{3}(t),\quad\text{with $u,t\in{\mathbb{C}}^{\times}$.}$
By our discussion in the previous subsection, the action of $s(u,t)\cdot\theta^{\vee}$ on $\mathfrak{spin}_{8}$ stabilizes $\mathfrak{t}^{\vee}={\rm Lie}(T^{\vee})$ and the root subspaces for the ${\rm G}_{2}$-root system. Hence,
$\mathfrak{spin}_{8}^{s(u,t)\cdot\theta^{\vee}}=(\mathfrak{t}^{\vee})^{\theta^{\vee}}\oplus\bigoplus_{\gamma\in\Phi_{G_{2}}}\mathfrak{u}_{\gamma}^{s(u,t)\cdot\theta^{\vee}}.$
In particular, to determine this Lie subalgebra fixed by $s(u,t)\cdot\theta^{\vee}$, we need to figure out its fixed space in each root subspace $\mathfrak{u}_{\gamma}$ for $\gamma\in\Phi_{G_{2}}$. This is a linear algebra computation which we illustrate for the case where $\gamma=\alpha$ and $\beta$.
  • $\gamma=\alpha$: the action of $s(u,t)\cdot\theta^{\vee}$ on the basis elements $\{x_{\alpha_{1}^{\vee}},x_{\alpha_{2}^{\vee}},x_{\alpha_{3}^{\vee}}\}$ is given by:
    $\begin{CD}x_{\alpha_{i}^{\vee}}@>{\theta^{\vee}}>{}>x_{\alpha_{i+1}^{\vee}}@>{{\rm Ad}(s(u,t))}>{}>t^{2}/u\cdot x_{\alpha_{i+1}^{\vee}}.\end{CD}$
    Hence, with respect to the basis $\{x_{\alpha_{i}^{\vee}}\}$, the action of $s(u,t)\cdot\theta^{\vee}$ on $\mathfrak{u}_{\alpha}$ is given by the matrix
    $t^{2}/u\cdot E=t^{2}/u\cdot\left(\begin{array}[]{ccc}0&0&1\\ 1&0&0\\ 0&1&0\end{array}\right).$
    The characteristic polynomial of $E$ is $\lambda^{3}+1$, so that its eigenvalues are the 3rd roots of unity. Hence, the 1-eigenspace of the $s(u,t)\cdot\theta^{\vee}$-action on $\mathfrak{u}_{\alpha}$ is nonzero if and only if $t^{6}=u^{3}$.
  • $\gamma=\beta$: the action of $s(u,t)\cdot\theta^{\vee}$ on $x_{\alpha_{0}^{\vee}}$ is given by
    $x_{\alpha_{0}^{\vee}}\mapsto u^{2}/t^{3}\cdot x_{\alpha_{0}^{\vee}}.$
    Hence $x_{\alpha_{0}^{\vee}}$ is fixed if and only if $u^{2}=t^{3}$.
We summarize the results of the computation for the other root spaces in the following table (in which we only indicate the results for positive roots).
Long roots $\theta$-orbit of 1-eigenspace nonzero iff
$\beta$ $\alpha_{0}^{\vee}$ $t^{3}=u^{2}$
$3\alpha+\beta$ $\alpha_{0}^{\vee}+\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\alpha_{3}^{\vee}$ $t^{3}=u$
$3\alpha+2\beta$ $2\alpha_{0}^{\vee}+\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\alpha_{3}^{\vee}$ $u=1$
Short roots $\theta$-orbit of 1-eigenspace nonzero iff
$\alpha$ $\alpha_{1}^{\vee}$ $t^{6}=u^{3}$
$\alpha+\beta$ $\alpha_{0}^{\vee}+\alpha_{1}^{\vee}$ $t^{3}=u^{3}$
$2\alpha+\beta$ $\alpha_{0}^{\vee}+\alpha_{1}^{\vee}+\alpha_{2}^{\vee}$ $t^{3}=1$
Using the above data, our job now is to find all pairs $(u,t)\in{\mathbb{C}}^{\times}\times{\mathbb{C}}^{\times}$ such that $\mathfrak{spin}_{8}^{s(u,t)\cdot\theta^{\vee}}$ is semisimple. Since we are interested in doing so up to conjugacy, it is helpful to first make a couple of elementary observations:
  • (a)
    If $\zeta$ is a 3rd root of unity, then $s(u,t)\cdot\theta^{\vee}$ and $s(u,t\zeta)\cdot\theta^{\vee}$ are conjugate by the element
    $\alpha_{1}(1)\cdot\alpha_{2}(\zeta)\cdot\alpha_{3}(\zeta^{2})\in T^{\vee}({\mathbb{C}}).$
  • (b)
    The elements $s(u,1)\cdot\theta^{\vee}$ and $s(u^{-1},1)\cdot\theta^{\vee}$ are conjugate by the Weyl group element $w_{\alpha_{0}}$ associated to the simple reflection in $\alpha_{0}$.
We can now begin our analysis:
  • (i)
    Referring to the 3 equations for the long root spaces (in the last column of the above table) as long root equations, we observe that if all 3 long root equations hold, then one has $u=1$ and $t^{3}=1$, so that the 3 short root equations also hold. By observation (a) above, we may take $t=1$ as well. Hence, one obtains $(u,t)=(1,1)$ and $\mathfrak{spin}_{8}^{\theta^{\vee}}=\mathfrak{g}_{2}$.
  • (ii)
    If two of the long root equations hold, then so does the last one, and hence we are reduced to the case above.
  • (iii)
    If exactly one of the long root equations holds, let us suppose without loss of generality that the last one holds, i.e. $u=1$, but $t^{3}\neq 1$. Then we observe that among the short root equations, only the first one has a chance of holding, and this happens precisely when $t^{6}=1$ but $t^{3}\neq 1$. If this short root equation does not hold, the fixed space is not semisimple and hence the resulting endoscopic group will not be elliptic. Hence, the only candidate for an elliptic endoscopic group here is for this equation to hold. By observation (a) above, there is no loss of generality in taking t=1t=-1, so that we have (u,t)=(1,1)(u,t)=(1,-1), and $\mathfrak{spin_{8}}^{s(1,-1)\cdot\theta^{\vee}}=\mathfrak{so}_{4}$.
  • (iv)
    Suppose that none of the long root equations hold. Then we need at least 2 of the short root equations to hold. But then all 3 of the short root equations will hold. Then we deduce that $u^{3}=1=t^{3}$. By (a) again, we may take $t=1$, and by (b), we see that $s(\zeta,1)\cdot\theta^{\vee}$ and $s(\zeta^{-1},1)\cdot\theta^{\vee}$ are conjugate. Hence we have $(u,t)=(\zeta,1)$ and $\mathfrak{spin}_{8}^{s(\zeta,1)\cdot\theta^{\vee}}=\mathfrak{sl_{3}}$.
3.2. Elliptic endoscopic data.
From the Lie algebra computations of the previous subsection, we can deduce:
Proposition 3.2.
The elliptic endoscopic groups for the twisted space $({\rm PGSO}_{8},\theta)$ determined by the triality automorphism are the split groups ${\rm G}_{2}$, ${\rm SL}_{3}$ and ${\rm SO}_{4}$. Moreover, the outer automorphism group of these elliptic endoscopic data is trivial.
Proof.
As mentionned above, because ${\rm Spin}_{8}({\mathbb{C}})$ is simply-connected, the group ${\rm Spin}_{8}({\mathbb{C}})^{s\cdot\theta^{\vee}}$ is in fact connected [KSh, Chap. 1, §1.1, Pg. 14]. The Lie algebra computations above thus allow us to identify ${\rm Spin}_{8}({\mathbb{C}})^{s\cdot\theta^{\vee}}$ up to isogeny, so we need to pin down the isogeny class. In item (i) above, we already know that:
${\rm Spin}_{8}({\mathbb{C}})^{\theta^{\vee}}={\rm G}_{2}({\mathbb{C}}).$
For item (iv), observe that $s(\zeta,1)\cdot\theta^{\vee}$ has order $3$ and thus provide a splitting of the automorphism short exact sequence (2.1) over ${\mathbb{Z}}/3{\mathbb{Z}}$. As we have noted, by [KMRT, Cor. 35.10 and Prop. 36.14], there are only two such splittings (up to conjugacy) with fixed groups of type ${\rm G}_{2}$ or ${\rm PGL}_{3}$. Hence, in the context of item (iv), we have
${\rm Spin}_{8}({\mathbb{C}})^{s(\zeta,1)\cdot\theta^{\vee}}\simeq{\rm PGL}_{3}({\mathbb{C}}).$
Finally, for item (iii), we observe that
$\mathfrak{spin_{8}}^{s(1,-1)\cdot\theta^{\vee}}\subset\mathfrak{spin}_{8}^{\theta^{\vee}}=\mathfrak{g}_{2}$
so that ${\rm Spin}_{8}({\mathbb{C}})^{s(1,-1)\cdot\theta^{\vee}}$ is a subgroup of ${\rm Spin}_{8}({\mathbb{C}})^{\theta^{\vee}}=G_{2}({\mathbb{C}})$ of type $A_{1}\times A_{1}$. This implies
${\rm Spin}_{8}({\mathbb{C}})^{s(1,-1)\cdot\theta^{\vee}}\simeq{\rm SO}_{4}({\mathbb{C}}).$
The first assertion of the proposition then follows from the definition of twisted endoscopic data given in [KSh, Chap. 2].
For the assertion on the triviality of automorphism groups, it suffices to show that for each of the three possibilities above for $H^{\vee}={\rm Spin}_{8}({\mathbb{C}})^{s\cdot\theta^{\vee}}$, the normalizer $N$ of $H^{\vee}$ in ${\rm Spin}_{8}({\mathbb{C}})$ is $H^{\vee}\times Z$ with $Z=Z_{{\rm Spin}_{8}}$.
Firstly, an inspection of the $8$-dimensional representation ${\rm std}_{|H^{\vee}}$ (which is respectively of type $1+7$, irreducible, or $1+3+4$), shows that the centralizer of $H^{\vee}$ in ${\rm Spin}_{8}({\mathbb{C}})$ is $Z_{H^{\vee}}\times Z$ in all cases. This shows the claim for the ${\rm G}_{2}$ case, since ${\rm Out}({\rm G}_{2})\,=1$. In the two other cases we have ${\rm Out}(\widehat{H})\simeq{\mathbb{Z}}/2$ and so some additional arguments are needed:
    (${\rm SO}_{4}$ case) We have $N=H^{\vee}$ since the $3$-dimensional summand of ${\rm std}_{|H^{\vee}}$ is not isomorphic to its outer conjugate.
    (${\rm PGL}_{3}$ case) Assume there is $g\in N$ such that ${\rm int}_{g}$ induces $h\mapsto{}^{\rm t}h^{-1}$ on $H^{\vee}\simeq{\rm PGL}_{3}$. Then $g$ commutes with ${\rm SO}_{3}({\mathbb{C}})=(H^{\vee})^{{\rm int}_{g}}$. Now ${\rm std}_{|{\rm SO}_{3}({\mathbb{C}})}$ has type $3+5$, and the element $g$ has to preserve each of these summands since it commutes with ${\rm SO}_{3}({\mathbb{C}})$ and thus acts by a scalar on each summand. Indeed, since ${\rm std}(g)$ belongs to $({\rm O}_{3}({\mathbb{C}})\times{\rm O}_{5}({\mathbb{C}}))\cap{\rm SO}_{8}({\mathbb{C}})$, the action of $g$ on the two summands must be via the same sign $\pm 1$. This implies that $g\in Z$, which is a contradiction.
Hence we have shown that $N=H^{\vee}\times Z$ in all cases, so that the outer automorphism group of each of these elliptic endoscopic data is trivial. ∎
3.3. Twisted Levi subspaces
We now describe the twisted Levi subspaces for the split $({\rm PGSO}_{8},\theta)$. These correspond to $\theta$-stable Levi subgroups of ${\rm PGSO}_{8}$. Up to conjugacy, these in turn correspond to the $\theta$-stable subsets of the set $\Delta(T,B)$ of simple roots. Hence we may enumerate the proper twisted Levi subspaces as follows:
  • the empty subset of $\Delta(T,B)$: this corresponds to the $\theta$-stable maximal torus $T$, giving rise to the twisted Levi subsapce $T\cdot\theta$.
  • the subset $\{\alpha_{0}\}$: this corresponds to a $\theta$-stable Levi subgroup $L_{0}\simeq({\rm GL}_{2}\times\mathbb{G}_{m}^{3})/\mathbb{G}_{m}^{\Delta}$, where the action of $\theta$ is given by cyclic permutation of the three $\mathbb{G}_{m}$’s.
  • the subset $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$: this corresponds to a $\theta$-stable Levi subgroup $L\simeq({\rm GL}_{2}^{3}\times\mathbb{G}_{m})/j(\mathbb{G}_{m}^{3})^{\Delta}$, where $j$ identifies $\mathbb{G}_{m}^{\Delta}$ with the center of ${\rm GL}_{2}^{3}$ and maps to $\mathbb{G}_{m}$ by the product map. The action of $\theta$ is given by cyclic permutation of the three ${\rm GL}_{2}$’s.
Now the theory of twisted endoscopy applies to each of the above twisted Levi subspaces as well. Hence one may determine their elliptic twisted endoscopic groups following a similar computation as in the previous subsection. In fact, from the general theory of twisted endoscopy, the elliptic twisted endoscopic groups of a twisted Levi subspace of $({\rm PGSO}_{8},\theta)$ are naturally identified with Levi subgroups of the elliptic twisted endoscopic groups $H$ of $({\rm PGSO}_{8},\theta)$. Since the groups $H$ are of rank $2$, besides their maximal torus $T_{H}$, their other proper Levi subgroups are the maximal ones and these are all isomorphic to ${\rm GL}_{2}$. More precisely:
  • For $H={\rm SL}_{3}$, there is a unique conjugacy class of maximal Levi subgroups, which we will denote by ${\rm GL}_{2,l}$;
  • For $H=G_{2}$, there are two conjugacy classes of maximal Levi subgroups, which we will denote by ${\rm GL}_{2,l}$ and ${\rm GL}_{2,s}$, corresponding to the long and short root ${\rm GL}_{2}$’s respectively;
  • For $H={\rm SO}_{4}$, there are two conjugacy classes of maximal Levi subgroups, which we will denote by ${\rm GL}_{2,l}$ and ${\rm GL}_{2,s}$.
The nomenclature is chosen so that the groups labelled as ${\rm GL}_{2,l}$ above are naturally identified with each other via admissible isomoprhisms, and likewise for the groups labelled as ${\rm GL}_{2,s}$. These admissible isomorphisms arise from the fact that $\xi_{H}({\rm GL}_{2,l}^{\vee})$ gives the same subgroup up to conjugacy in ${\rm Spin}_{8}({\mathbb{C}})$ for the various $H$. Likewise the various $\xi_{H}({\rm GL}_{2,s}^{\vee})$ are the same subgroups in ${\rm Spin}_{8}({\mathbb{C}})$ up to conjugacy.
With the above notations, we can now tabulate the elliptic twisted endoscopic groups of the twisted Levi subspaces, as well as the Levi subgroups of $H$ they are identified with. In the table, we have written
$^{1-\theta}T:=\text{ image of $[(1-\theta):T\rightarrow T]$.}$
Twisted Levi subspace $(T,\theta)$ $(L_{0},\theta)$ $(L,\theta)$
Elliptic twisted endoscopic group $T/^{1-\theta}T$ ${\rm GL}_{2}$ ${\rm GL}_{2}$
Levi subgroup of $H$ identified with $T_{H}$ ${\rm GL}_{2,s}$ ${\rm GL}_{2,l}$
4. Local Twisted Endoscopic Transfer
Let $F$ be a local field of characteristic 0. Having determined the twisted endoscopic groups $H$ for $({\rm PGSO}_{8},\theta)$, the theory of twisted endoscopy furnishes a local transfer map from $\theta$-twisted orbital integrals on ${\rm PGSO}_{8}(F)$ to orbital integrals on $H(F)$. In this section, we recall the basic objects and record some results on this local transfer that we will need.
4.1. Local K-forms
A slight complication with the theory of stable twisted trace formula is the need to take into account of K-forms [MW1, §I.1.11], which are certain inner forms of the twisted space in question.
In the context of ${\rm PGSO}_{8}$, we have seen in §2 that an octonion algebra $\mathbb{O}$ gives rise to a group ${\rm PGSO}_{8}^{\mathbb{O}}$, which is an inner form of the corresponding split group equipped with a triality automorphism. Over a number field or a local field, these turn out to be the K-forms of the twisted space $({\rm PGSO}_{8},\theta)$. Thus, over a non-Archimedean local field or $\mathbb{C}$}, $({\rm PGSO}_{8},\theta)$ has no additional $K$-forms, since there is a unique octonion algebra up to isomorphism. On the other hand, over $\mathbb{R}$}, there are two octonions algebras: the split one and a division algebra. As such, there are two K-forms of $({\rm PGSO}_{8},\theta)$ over $\mathbb{R}$}.
4.2. Twisted orbital integrals
We now recall some properties of the space $\mathcal{I}^{\theta}({\rm PGSO}^{\mathbb{O}}_{8})$ of $\theta$-twisted orbital integrals on a local K-form ${\rm PGSO}^{\mathbb{O}}_{8}$ over $F$. For simplicity, we will suppress $\mathbb{O}$ from the notation and write $G={\rm PGSO}_{8}$.
For a test function $f\in C^{\infty}_{c}(G(F))$, its normalized $\theta$-twisted orbital integral $\mathcal{O}^{\theta}(-,f)$ is a function on the subset of $\theta$-regular semisimple elements $\gamma$ of $G(F)$ defined by
(4.1) $\mathcal{O}^{\theta}(\gamma,f)=\Delta_{{\rm PGSO}_{8}}(\gamma\cdot\theta)\cdot\int_{G_{\gamma\cdot\theta}(F)\backslash G(F)}f(g^{-1}\gamma\theta(g))\,\frac{dg}{dt},$
where $G_{\gamma\cdot\theta}$ denotes the stabilizer of $\gamma\cdot\theta$ in $G$ and
$\Delta_{{\rm PGSO}_{8}}(\gamma\cdot\theta)=|\det(1-{\rm Ad}(t\theta)|\mathfrak{g}/\mathfrak{g}_{\gamma\cdot\theta})|^{1/2}$
is the positive square root of the $\theta$-twisted Weyl discriminant. The space $\mathcal{I}^{\theta}(G)$ is the topological vector space of such normalized twisted orbital integrals.
Assuming now that $G$ is split, we have the following structures on the space $\mathcal{I}^{\theta}(G)$:
  • (i)
    [MW1, I.3.1]   For each $\theta$-stable Levi subgroup $M$ of $G={\rm PGSO}_{8}$ (or equivalently any twisted Levi subspace $M\cdot\theta$), there is a restriction or “constant term” map:
    ${\rm rest}^{G\cdot\theta}_{M\cdot\theta}:\mathcal{I}^{\theta}(G)\longrightarrow\mathcal{I}^{\theta}(M)^{W(G,M)^{\theta}}$
    where $W(G,M)=N_{G(F)}(M)/M(F)$ and $W(G,M)^{\theta}$ is the $\theta$-fixed subgroup. Moreover, one has the transitivity of restrictions:
    ${\rm rest}^{M\cdot\theta}_{T\cdot\theta}\circ{\rm rest}^{G\cdot\theta}_{M\cdot\theta}={\rm rest}^{G\cdot\theta}_{T\cdot\theta}$
    where $M=L_{0}$ or $L$.
  • (ii)
    [MW1, I.4.2]  The restriction maps give rise to a natural filtration
    $0=\mathcal{F}^{-1}\subset\mathcal{F}^{0}\subset\mathcal{F}^{1}\subset\mathcal{F}^{2}=\mathcal{I}^{\theta}({\rm PGSO}_{8}).$
    These subspaces are given by:
    $\mathcal{F}^{1}={\rm Ker}({\rm rest}^{G\cdot\theta}_{T\cdot\theta})\quad\text{and}\quad\mathcal{F}^{0}={\rm Ker}({\rm rest}^{G\cdot\theta}_{L_{0}\cdot\theta})\cap{\rm Ker}({\rm rest}^{G\cdot\theta}_{L\cdot\theta}).$
    The subspace $\mathcal{F}^{0}$ is also called the cuspidal subspace and is denoted by $\mathcal{I}^{\theta}_{cusp}(G)$.
  • (iii)
    [MW1, I.4.2]  The successive quotients of the above filtration are given by
    $\mathcal{F}^{2}/\mathcal{F}^{1}\simeq\mathcal{I}^{\theta}_{cusp}(T)^{W(G,T)^{\theta}}$
    and
    $\mathcal{F}^{1}/\mathcal{F}_{0}\simeq\mathcal{I}^{\theta}_{cusp}(L_{0})^{W(G,L_{0})^{\theta}}\oplus\mathcal{I}^{\theta}_{cusp}(L)^{W(G,L)^{\theta}}.$
    The isomorphisms are naturally induced by the restriction maps.
Remark: In (i) above, we have used the group $W(G,M)^{\theta}$, whereas in [MW1, I.3.1 and I.4.2], one sees instead
$W(M\cdot\theta)=N_{G(F)}(M\cdot\theta)/M(F).$
Let us show that these are the same. Suppose that $g\in W(M\cdot\theta)$, so that $g^{-1}M\theta(g)=M$. This implies that $gMg^{-1}=M\cdot\theta(g)g^{-1}$. For this to hold, one must have $\theta(g)g^{-1}\in M$, so that $g$ normalizes $M$ and its image in $W(G,M)=N_{G(F)}(M)/M(F)$ is fixed by $\theta$. Conversely, if $g\in N_{G(F)}(M)$ is such that $\theta(g)g^{-1}\in M(F)$, then
$g^{-1}M\cdot\theta\cdot g=g^{-1}M\cdot\theta(g)\cdot\theta=g^{-1}Mg\cdot\theta=M\cdot\theta,$
so that $g\in N_{G(F)}(M\cdot\theta)$.
4.3. Stable orbital integrals
For an connected reductive group $H$ over $F$, let $\mathcal{SI}(H)$ denote the space of stable orbital integrals on $H(F)$. Recall that for a test function $f_{H}\in C^{\infty}_{c}(H(F))$, its normalized orbital integral is a function on strongly regular semisimple elements $\gamma$ of $H(F)$ defined by
[…]
where $H_{\gamma}$ is the stabilizer of $\gamma$ in $H(F)$ (and hence is a maximal $F$-torus) and
$\Delta_{H}(\gamma)=|\det(1-{\rm Ad}(\gamma)|\mathfrak{h}/\mathfrak{h}_{\gamma}))|^{1/2}$
is the positive square root of the Weyl discriminant. We may further replace the normalized orbital integral above by its stabilized version:
(4.3) $\mathcal{SO}(\gamma,f_{H})=\sum_{\gamma^{\prime}}\mathcal{O}(\gamma^{\prime},f_{H})$
with the sum running over the conjugacy classes in $H(F)$ which are stably conjugate to $\gamma$. The space spanned by such normalized stable orbital integrals is the topological vector space $\mathcal{SI}(H)$.
We have the analogous structures on $\mathcal{SI}(H)$:
  • (i)
    For each Levi subgroup $M\subset H$, there is a restriction or “constant term” map:
    ${\rm rest}_{M}^{H}:\mathcal{SI}(H)\longrightarrow\mathcal{SI}(M)^{W(H,M)}$
    where $W(H,M)=N_{H(F)}(M)/M(F)$.
  • (ii)
    The restriction maps give rise to a natural filtration on $\mathcal{SI}(H)$:
    $0=:\mathcal{F}^{-1}\subset\mathcal{F}^{0}\subset\mathcal{F}^{1}\subset...\subset\mathcal{F}^{r}=\mathcal{SI}(H).$
    The subspace $\mathcal{F}^{k}$ is defined by
    $\mathcal{F}^{k}=\bigcap_{M\in\mathcal{L}_{>k}(H)}{\rm Ker}({\rm rest}_{M}^{H})$
    where the intersection runs over the set $\mathcal{L}_{>k}(H)$ of representatives of $H(F)$-conjugacy classes of Levi subgroups $M$ of $H$ with ${\rm corank}(M):=\dim Z_{M}/Z_{H}>k$ (here $Z_{H}$ and $Z_{M}$ denote the centers of $H$ and $M$). In particular, $\mathcal{F}^{0}$ denotes the subspace of elements whose restriction to any proper Levi subgroup vanish. This subspace is also called the cuspidal subspace and is denoted by $\mathcal{SI}_{cusp}(H)$.
  • (iii)
    The successive quotients of the above filtration is given by:
    $\mathcal{F}^{k}/\mathcal{F}^{k-1}\simeq\bigoplus_{M\in\mathcal{L}_{k}(H)}\mathcal{SI}_{cusp}(M)^{W(H,M)}$
    where the sum runs over a set $\mathcal{L}_{k}(H)$ of representatives of $H(F)$-conjugacy classes of Levi subgroups $M$ with corank($M$) =k=k. The isomorphism is given by the restriction maps to Levi subgroups of corank $k$.
We will of course be applying the above results to the case when $H$ is an elliptic twisted endoscopic group of $({\rm PGSO}_{8},\theta)$. As an illustration, let us explicate the above general results when $H={\rm SL}_{3}$. In this case, using the notations from §3.3, we have
$\mathcal{L}_{1}(H)=\{{\rm GL}_{2,l}\}\quad\text{and}\quad\mathcal{L}_{2}(H)=\{T_{{\rm SL}_{3}}\}.$
Then the above results say that
(4.4) $\mathcal{F}^{1}/\mathcal{F}^{0}\simeq\mathcal{SI}_{cusp}({\rm GL}_{2,l})=\mathcal{I}_{cusp}({\rm GL}_{2,l})$
and
(4.5) $\mathcal{F}^{2}/\mathcal{F}^{1}\simeq\mathcal{SI}(T_{{\rm SL}_{3}})^{W({\rm SL}_{3},T_{{\rm SL}_{3}})}=\mathcal{I}(T_{{\rm SL}_{3}})^{S_{3}}.$
Here, for the first identity, we note that the Weyl group $W({\rm SL}_{3},{\rm GL}_{2,l})$ is trivial and any invariant distribution on ${\rm GL}_{2,l}$ is already stably-invariant, whereas for the second identity, we have $W({\rm SL}_{3},T_{{\rm SL}_{3}})\simeq S_{3}$.
4.4. Local transfer
For each local $K$-form $({\rm PGSO}_{8}^{\mathbb{O}},\theta)$ and each twisted endoscopic group $H$, the theory of twisted endoscopy furnishes a local transfer of orbital integrals from $({\rm PGSO}_{8}^{\mathbb{O}},\theta)$ to $H$. More precisely, we have:
Proposition 4.6.
Let $F$ be a local field of characteristic 0 and consider the setting of $({\rm PGSO}_{8},\theta)$ over $F$. Then we have:
(i) The twisted endoscopic transfer gives an isomorphism of topological vector spaces
${\rm trans}:\bigoplus_{\mathbb{O}}\mathcal{I}^{\theta}_{cusp}({\rm PGSO}^{\mathbb{O}}_{8})\longrightarrow\mathcal{SI}_{cusp}({\rm G}_{2})\oplus\mathcal{SI}_{cusp}({\rm SO}_{4})\oplus\mathcal{SI}_{cusp}({\rm SL}_{3}),$
where the sum runs over the isomorphism classes of octonion $F$-algebras.
(ii) More generally, the twisted endoscopic transfer gives an injective map
${\rm trans}:\bigoplus_{\mathbb{O}}\mathcal{I}^{\theta}({\rm PGSO}^{\mathbb{O}}_{8})\longrightarrow\mathcal{SI}({\rm G}_{2})\oplus\mathcal{SI}({\rm SO}_{4})\oplus\mathcal{SI}({\rm SL}_{3}).$
Its image is the subspace of “compatible families” of stably invariant distributions of its target (a notion we will recall after the proposition; see [MW1, I.4.11]).
(iii) Let $F$ be a non-Archimedean local field whose residue characteristic is sufficiently large. For each of the group $G$ under consideration here, let $\mathcal{H}_{G}$ denote the spherical Hecke algebra of $G(F)$ with respect to a hyperspecial subgroup. For each elliptic twisted endoscopic group $H$ of $({\rm PGSO}_{8},\theta)$, the embedding $\xi_{H}:H^{\vee}\hookrightarrow{\rm Spin}_{8}({\mathbb{C}})$ induces (via the Satake isomorphism) a natural algebra homomorphism
$\xi_{H}^{*}:\mathcal{H}_{{\rm PGSO}_{8}}\longrightarrow\mathcal{H}_{H}.$
Then one has a commutative diagram (the fundamental lemma for spherical Hecke algebras):
[…]
where the vertical arrows are the formation of $\theta$-twisted orbital integrals and stable orbital integrals respectively.
Proof.
For (i) and (ii), see [MW1, Prop. I.4.11, Pg. 96]. For the fundamental lemma in (iii), see [MW1, §I.6.4, Pg. 165] and the references therein. ∎
We have still to define the notion of a “compatible family” of stably invariant distributions in Proposition 4.6(ii) above. Suppose that
$(f_{G_{2}},f_{{\rm SO}_{4}},f_{{\rm SL}_{3}})\in\mathcal{SI}({\rm G}_{2})\oplus\mathcal{SI}({\rm SO}_{4})\oplus\mathcal{SI}({\rm SL}_{3}).$
We say that this is a compatible family if
${\rm rest}^{G_{2}}_{{\rm GL}_{2,l}}(f_{G_{2}})={\rm rest}^{{\rm SO}_{4}}_{{\rm GL}_{2,l}}(f_{{\rm SO}_{4}})={\rm rest}^{{\rm SL}_{3}}_{{\rm GL}_{2,l}}(f_{{\rm SL}_{3}}),$
and
${\rm rest}^{G_{2}}_{{\rm GL}_{2,s}}(f_{G_{2}})={\rm rest}^{{\rm SO}_{4}}_{{\rm GL}_{2,s}}(f_{{\rm SO}_{4}}).$
Here, we recall that the Levi subgroups of the various twisted endoscopic groups $H$ labelled as ${\rm GL}_{2,l}$ (respectively ${\rm GL}_{2,s}$) are naturally identified with each other via admissible isomorphisms, so that the above identities make sense. The transitivity of restriction maps implies that when the above two identities hold, we also have
${\rm rest}^{G_{2}}_{T_{G_{2}}}(f_{G_{2}})={\rm rest}^{{\rm SO}_{4}}_{T_{{\rm SO}_{4}}}(f_{{\rm SO}_{4}})={\rm rest}^{{\rm SL}_{3}}_{T_{{\rm SL}_{3}}}(f_{{\rm SL}_{3}}).$
For any elliptic twisted endoscopic group $H=G_{2}$, ${\rm SL}_{3}$ or ${\rm SO}_{4}$ of $({\rm PGSO}_{8},\theta)$, the projection of the map $\rm trans$} in (ii) to the $H$-factor gives
${\rm trans}_{H}:\bigoplus_{\mathbb{O}}\mathcal{I}^{\theta}({\rm PGSO}^{\mathbb{O}}_{8})\longrightarrow\mathcal{SI}(H).$
By Proposition 4.6(ii), the image of ${\rm trans}_{H}$ is precisely the subspace of those elements which can be completed to a compatible family. Moreover, the natural filtration $(\mathcal{F}^{\bullet})$ on $\mathcal{SI}(H)$ induces the filtration
$\mathcal{T}^{\bullet}:=\mathcal{F}^{\bullet}\cap{\rm trans}_{H}(I^{\theta}(G))\quad\text{ on ${\rm trans}_{H}(I^{\theta}(G))$,}$
so that
$\mathcal{T}^{k}/\mathcal{T}^{k-1}\subset\mathcal{F}^{k}/\mathcal{F}^{k-1}.$
The following proposition characterizes ${\rm trans}_{H}(I^{\theta}(G))$ in terms of this filtration and the associated graded pieces:
Proposition 4.7.
(i) Let $H=G_{2}$. Then the successive quotients of the filtration $(\mathcal{T}^{\bullet})$ are given by:
$\begin{cases}\mathcal{T}^{2}/\mathcal{T}^{1}=\mathcal{I}(T_{G_{2}})^{W(G_{2},T_{G_{2}})},\\ \mathcal{T}^{1}/\mathcal{T}^{0}=\mathcal{I}_{cusp}({\rm GL}_{2,l})^{{\rm Out}({\rm GL}_{2})}\oplus\mathcal{I}_{cusp}({\rm GL}_{2,s})^{{\rm Out}({\rm GL}_{2})},\\ \mathcal{T}^{0}=\mathcal{F}^{0}=\mathcal{I}_{cusp}(G_{2}).\end{cases}$
Here ${\rm Out}({\rm GL}_{2})$ refers to the outer automorphism group of ${\rm GL}_{2}$.
(ii) Let $H={\rm SL}_{3}$. Then the successive quotients of the filtration $(\mathcal{T}^{\bullet})$ are given by:
$\begin{cases}\mathcal{T}^{2}/\mathcal{T}^{1}=\mathcal{I}(T)^{W(G_{2},T_{G_{2}})},\\ \mathcal{T}^{1}/\mathcal{T}^{0}=\mathcal{I}_{cusp}({\rm GL}_{2,l})^{{\rm Out}({\rm GL}_{2})},\\ \mathcal{T}^{0}=\mathcal{F}^{0}=\mathcal{SI}_{cusp}({\rm SL}_{3}).\end{cases}$
Here, we have made use of the natural identification $T_{{\rm SL}_{3}}\simeq T_{G_{2}}$ to transport the action of $W(G_{2},T_{G_{2}})$ to $T_{{\rm SL}_{3}}$, and under this transport of structure, $W(G_{2},T_{G_{2}})=W({\rm SL}_{3},T_{{\rm SL}_{3}})\times S_{2}$ with the nontrivial element of $S_{2}$ acting by inverting on $T_{{\rm SL}_{3}}$.
(iii) Let $H={\rm SO}_{4}$. Then the successive quotients of the filtration $(\mathcal{T}^{\bullet})$ are given by:
$\begin{cases}\mathcal{T}^{2}/\mathcal{T}^{1}=\mathcal{I}(T)^{W(G_{2},T_{G_{2}})},\\ \mathcal{T}^{1}/\mathcal{T}^{0}=\mathcal{I}_{cusp}({\rm GL}_{2,l})^{{\rm Out}({\rm GL}_{2})}\oplus\mathcal{I}_{cusp}({\rm GL}_{2,s})^{{\rm Out}({\rm GL}_{2})},\\ \mathcal{T}^{0}=\mathcal{F}^{0}=\mathcal{I}_{cusp}({\rm SO}_{4}).\end{cases}$
Here, we have again made use of the natural identification $T_{{\rm SO}_{4}}\simeq T_{G_{2}}$ to transport the action of $W(G_{2},T_{G_{2}})$ to $T_{{\rm SO}_{4}}$.
Proof.
This can be deduced from [MW1, §I.4.3, Lemme, Pg. 80] and the definition of “compatible families”; we omit the details. ∎
Corollary 4.8.
The map ${\rm trans}_{H}$ is surjective if $H=G_{2}$, but is not surjective if $H={\rm SL}_{3}$ or ${\rm SO}_{4}$. When $H={\rm SL}_{3}$, the image of ${\rm trans}_{{\rm SL}_{3}}$ contains the subspace $\mathcal{SI}({\rm SL}_{3})^{{\rm Out}({\rm SL}_{3})}$ of $\mathcal{SI}({\rm SL}_{3})$ fixed by the outer automorphism group of ${\rm SL}_{3}$.
Proof.
For the case $H={\rm SL}_{3}$, one may compare Proposition 4.7(i) and the equations (4.4) and (4.5). The other cases are similar and so we omit the details. ∎
4.5. Spectral transfer
Let us now specialize to the case $H={\rm SL}_{3}$. The dual of the transfer map ${\rm trans}_{{\rm SL}_{3}}$ is a map
${\rm trans}_{{\rm SL}_{3}}^{*}:\mathcal{SI}({\rm SL}_{3})^{*}\longrightarrow\mathcal{I}({\rm PGSO}_{8})^{*}$
from the space of stably invariant distributions of $SL_{3}(F)$ to the space of invariant distributions on ${\rm PGSO}_{8}(F)$. The fact that the transfer map ${\rm trans}_{{\rm SL}_{3}}$ is not surjecitive implies that this dual transfer map is not injective. However, we note:
Lemma 4.9.
Suppose that $\pi$ is a semisimple admissible representation of ${\rm SL}_{3}(F)$ whose Harish-Chandra character $\Theta_{\pi}$ is a stable distribution. Then ${\rm trans}_{{\rm SL}_{3}}^{*}(\Theta_{\pi})\neq 0$.
Proof.
By the description of the image of ${\rm trans}_{{\rm SL}_{3}}$ given in Corollary 4.8, it suffices to show that $\Theta_{\pi}$ is not identically zero on $\mathcal{SI}({\rm SL}_{3})^{{\rm Out}({\rm SL}_{3})}$. Let $\sigma$ be a nontrivial outer automorphism of ${\rm SL}_{3}$. Then the space $\mathcal{SI}({\rm SL}_{3})^{{\rm Out}({\rm SL}_{3})}$ is spanned by the stable orbital integrals of test functions of the form $f+f^{\sigma}$. Now since $\pi^{\sigma}\simeq\pi^{\vee}$, we have
$\Theta_{\pi}(f^{\sigma})=\Theta_{\pi^{\sigma}}(f)=\Theta_{\pi^{\vee}}(f),$
so that
$\Theta_{\pi}(f+f^{\sigma})=\Theta_{\pi\oplus\pi^{\vee}}(f).$
Since $\pi\oplus\pi^{\vee}$ is a nonzero admissible semisimple representation, it is thus clear that $\Theta_{\pi\oplus\pi^{\vee}}$ is nonzero on some test function $f$. ∎
5. Unramified Spectral Transfer
We now specialize to the setting of a non-Archimedean local field $F$ with ring of integers $\mathcal{O}_{F}$ (so that there is a unique K-form, namely the split ${\rm PGSO}_{8}$). Recall that in §2.4, we have started with the Chevalley group ${\rm PGSO}_{8}$ defined over $\mathbb{Z}$} with a given Chevalley-Steinberg system of épinglage and equipped with a triality automorphism $\theta$. Hence we have a hyperspecial subgroup $K=G(\mathcal{O}_{F})$ which is stablized by $\theta$. Since ${\rm PGSO}_{8}$ is adjoint, there is a unique conjugacy class of hyperspecial subgroups.
5.1. Unramified L-packets
Now let $H$ be one of the elliptic endoscopic groups of $({\rm PGSO}_{8},\theta)$. In fact, we shall only be interested in the case $H={\rm SL}_{3}$ and the reader can assume this is the case in the following. Suppose that we have an unramified L-packet of $H$ associated to a semisimple conjugacy class $s\in H^{\vee}$. Recall that $s$ determines an unramified character $\chi_{s}$ of a maximal split torus $T_{H}$ of $H$ and the L-packet in question consists simply of the constituents of the corresponding principal series representation $I_{H}(\chi_{s})$ (parabolically induced from $\chi_{s}$) which have nonzero invariant vectors under some hyperspecial subgroups of $H(F)$. Let us assume for simplicity that the unramified constituents of $I_{H}(\chi_{s})$ are generic and unitary. In that case, $I_{H}(\chi_{s})$ is semisimple and the L-packet consists of all its irreducible constituents.
The stable trace $\Theta_{s}$ of such an unramified L-packet is the sum of the Harish-Chandra character of its members and hence (under our hypothesis above) is simply the Harish-Chandra character of $I_{H}(\chi_{s})$. It is a stable distribution on $H(F)$ and hence defines a linear functional on $\mathcal{SI}(H)$. We may consider the pullback of $\Theta_{s}$ under the transfer map ${\rm trans}_{H}$ of Proposition 4.6(iii). For the purpose of applying the trace formula, we need to determine the unramified spectral transfer $\Theta_{s}\circ{\rm trans}_{H}$ precisely.
To formulate the answer, consider the unramified L-packet of ${\rm PGSO}_{8}(F)$ corresponding to the semisimple conjugacy class of $\xi_{H}(s)$, where $\xi_{H}:H^{\vee}\hookrightarrow{\rm Spin}_{8}({\mathbb{C}})$ is the natural embedding. This unramified L-packet consists of the unramified constituents of an unramified principal series representation which can be described as follows. The restriction of the embedding $\xi_{H}$ defines an isomorphism
$\xi_{H}:T_{H}^{\vee}\longrightarrow(T^{\vee})^{\theta^{\vee}}\subset T^{\vee},$
where $T$ is a maximal split torus of ${\rm PGSO}_{8}$. This injective map of dual torus arises from a surjective morphism
(5.1) $\eta:T\longrightarrow T_{H}$
satisfying
(5.2) ${\rm Ker}(\eta)={}^{1-\theta}T:=\text{ image of $[(1-\theta):T\rightarrow T]$.}$
By pulling back via $\eta$, one obtains an unramified character $\chi_{s}\circ\eta$ of $T(F)$. Then the unramified principal series representation of ${\rm PGSO}_{8}(\mathbb{F})$ corresponding to $\xi_{H}(s)$ is the representation $I(\chi_{s}\circ\eta)$ parabolically induced from $\chi_{s}\circ\eta$.
5.2. Extensions to twisted space
Since $\xi_{H}(s)$ is fixed by $\theta^{\vee}$, the character $\chi_{s}\circ\eta$ is $\theta$-invariant, which implies that the isomorphism class of $I(\chi_{s}\circ\eta)$ is fixed by $\theta$ and hence can be extended to the semi-direct product ${\rm PGSO}_{8}(F)\rtimes\langle\theta\rangle$ (or rather to the twisted space ${\rm PGSO}_{8}(F)\cdot\theta$). The extension is however not unique. The following lemma gives several equivalent ways of fixing such an extension.
Lemma 5.3.
. The following three extensions of $I(\chi_{s}\circ\eta)$ to ${\rm PGSO}_{8}(F)\rtimes\langle\theta\rangle$ are the same:
  • (i)
    (Extension by induction) Extend $\chi_{s}\circ\eta$ to $T(F)\rtimes\langle\theta\rangle$ by letting $\theta$ act trivially; by the functoriality of induction, this gives rise to an action of $\theta$ on $I(\chi_{s}\circ\eta)$, described explicitly by
    $(\theta\cdot\phi)(g)=\phi(\theta^{-1}(g))$
    for any $\phi\in I(\chi_{s}\circ\eta)$ and $g\in{\rm PGSO}_{8}(F)$.
  • (ii)
    (Spherical-normalized extension) Extend $I(\chi_{s}\circ\eta)$ by requiring that $\theta$ acts as identity on the 1-dimensional space $I(\chi_{s}\circ\eta)^{K}$.
  • (iii)
    (Whittaker-normalized extension) For a nontrivial additive character $\psi$ of $F$ of consuctor $\mathcal{O}_{F}$, consider the Whittaker datum determined by $\psi$ and the fixed Chavalley-Steinberg system of épinglage and a corresponding nontrivial Whittaker functional $\mathcal{W}:I(\chi_{s}\circ\eta)\longrightarrow{\mathbb{C}}$. Then the Whittaker datum is fixed by $\theta$ and we extend $I(\chi_{s}\circ\eta)$ by requiring that
    $\mathcal{W}\circ\theta=\mathcal{W}.$
Proof.
Let $\phi_{0}\in I(\chi_{s}\circ\eta)$ be a nonzero $K$-unramfiied vector which is unique up to scaling. Then $\phi_{0}$ is characterized by the requirement that $\phi_{0}|_{K}$ is a constant function. Since $\theta$ stabilizes $K$, one deduces easily that the extensions in (i) and (ii) are the same. On the other hand, by the Casselman-Shalika formula, one knows that $\mathcal{W}(\phi_{0})\neq 0$. Hence the Whittaker-normalzied extension is also spherical-normalized, showing that the extensions in (ii) and (iii) are the same. ∎
We shall adopt the extension of $I(\chi_{s}\circ\eta)$ to ${\rm PGSO}_{8}(F)\rtimes\langle\theta\rangle$ supplied by the above lemma, denoting the resulting action of $\theta$ by $I(\theta)$. Using this extension, we define the $\theta$-twisted trace of $I(\chi_{s}\circ\eta)$ as the Harish-Chandra character of the extended representation, restricted to the coset ${\rm PGSO}_{8}(F)\cdot\theta$. More precisely, it is given by the linear functional
$\Theta^{\theta}_{\xi_{H}(s)}:f\mapsto{\rm Tr}(I(\theta)\circ I(\chi_{s}\circ\eta)(f)),\quad\text{ for $f\in C^{\infty}_{c}({\rm PGSO}_{8}(F))$.}$
which is $\theta$-conjugacy-invariant and hence factors to a linear functional on $\mathcal{I}^{\theta}({\rm PGSO}_{8})$. Observe that if $f\in\mathcal{H}_{{\rm PGSO}_{8}}$, then
${\rm Tr}(I(\theta)\circ I(\chi_{s}\circ\eta)(f))={\rm Tr}(I(\chi_{s}\circ\eta)(f)),$
since $I(\chi_{s}\circ\eta)(f)$ has image contained in $I(\chi_{s}\circ\eta)^{K}$ and $I(\theta)$ acts as identity on $I(\chi_{s}\circ\eta)^{K}$. Hence,
$\Theta^{\theta}_{\xi_{H}(s)}=\Theta_{\xi_{H}(s)}\quad\text{on $\mathcal{H}_{{\rm PGSO}_{8}}$.}$
5.3. Unramified spectral transfer
Now one has:
Proposition 5.4.
In the above context, one has the twisted endoscopic character identity:
$\Theta_{s}\circ{\rm trans}_{H}=\Theta^{\theta}_{\xi_{H}(s)}$
as linear functionals on $\mathcal{I}^{\theta}({\rm PGSO}_{8})$. In particular, the twisted endoscopic character identity defines a lifting of unramified L-packets from $H$ to ${\rm PGSO}_{8}$ which is functorial with respect to the embedding
$\xi_{H}:H^{\vee}\longrightarrow{\rm PGSO}_{8}^{\vee}={\rm Spin}_{8}({\mathbb{C}}).$
Proof.
For a test function $f\in C^{\infty}_{c}({\rm PGSO}_{8}(F))$, we need to show that
$\Theta_{s}({\rm trans}_{H}(f))=\Theta^{\theta}_{\xi_{H}(s)}(f).$
If $f\in\mathcal{H}_{{\rm PGSO}_{8}}$, then this identity follows readily from Proposition 4.6(iii). In the general case, the proof of this identity is similar to that of [CG, Prop. 6.2].
We first compute each side of the purported identity. For a test function $f_{H}\in C^{\infty}_{c}(H(F))$, we first compute $\Theta_{s}(f_{H})$. Recall that the Harish-Chandra character $\Theta_{s}$ of $I_{H}(\chi_{s})$ is given by a locally constant and locally integrable conjugacy-invariant function on the open dense set $H(F)_{rs}$ of regular semisiple elements of $H(F)$. We denote this function by $\Theta_{s}$ as well, so that
$\Theta_{s}(f_{H})=\int_{H(F)_{rs}}\Theta_{s}(x)\cdot f_{H}(x)\,dx.$
Moreover, since $\Theta_{s}$ is the character of a principal series representation, the function $\Theta_{s}$ is supported on the conjugacy classes of regular elements of the maximal split torus $T_{H}$. Hence, by the Weyl integration formula, one has (cf. [CG, §5.8])
$\Theta_{s}(f_{H})=\int_{T_{H}(F)_{rs}}\chi_{s}(t_{H})\cdot\mathcal{O}(t_{H},f_{H})\,dt_{H}$
where $\mathcal{O}(-,f_{H})$ is the normalized orbital integral of $f_{H}$ defined in (4.2). Since $t_{H}$ is a regular element in the maximal split torus, there is a unique conjugacy class in the stable conjugacy class of $t_{H}$, and hence one has $\mathcal{SO}(t_{H},f_{H})=\mathcal{O}(t_{H},f_{H})$, where $\mathcal{SO}(t_{H},f_{H})$ is the stable orbital integral defined in (4.3). Hence, we have:
(5.5) $\Theta_{s}(f_{H})=\int_{T_{H}(F)_{rs}}\chi_{s}(t_{H})\cdot\mathcal{SO}(t_{H},f_{H})\,dt_{H}.$
For a test function $f\in C^{\infty}_{c}({\rm PGSO}_{8}(F))$, we may likewise compute the twisted trace $\Theta^{\theta}_{\xi_{H}(s)}(f)$ analogously, using a twisted Weyl integration formula (see [CG, §5.7 and §5.8]). This gives
(5.6) $\Theta^{\theta}_{\xi_{H}(s)}(f)=\int_{{}^{1-\theta}T(F)\backslash T(F)_{\theta-rs}}\chi_{\xi_{H}(s)}(t)\cdot\mathcal{O}^{\theta}(t,f)\,dt$
where $T(F)_{\theta-rs}$ denotes the open dense subset of $\theta$-regular elements and $\mathcal{O}_{\theta}(t,f)$ is the $\theta$-twisted orbital integral of $f$ defined in (4.1). By (5.1) and (5.2), one has
$\eta:T/{}^{1-\theta}T\simeq T_{H}\quad\text{and}\quad\chi_{\xi_{H}(s)}=\chi_{s}\circ\eta.$
In addition, under the orbit correspondence in the theory of twisted endoscopy, the pair $(\eta(t),t)\in T_{H}(F)_{rs}\times T(F)_{\theta_{r}s}$ correspond. Moreover, the value of Whittaker-normalized twisted transfer factor on $(\eta(t),t)$ is given by
$\Delta(\eta(t),t)=1,$
so that by the definition of transfer,
$\mathcal{SO}(\eta(t),{\rm trans}_{H}(f))=\mathcal{O}^{\theta}(t,f).$
Hence, by a change of variable, setting $t_{H}=\eta(t)$,
(5.7) $\Theta^{\theta}_{\xi_{H}(s)}(f)=\int_{T_{H}(F)}\chi_{s}(t_{H})\cdot\mathcal{SO}(t_{H},{\rm trans}_{H}(f))\,dt_{H}.$
Comparing (5.5) and (5.7) and taking $f_{H}={\rm trans}_{H}(f)$ in (5.5), we deduce the desired unramified spectral transfer identity. ∎
It seems to us that this unramified spectral transfer could and should be shown in the general context of twisted endoscopy; this will eliminate the need to deal with it on a case-by-case basis.
5.4. The case $H={\rm SL}_{3}$
Let us specialize further to the case of the twisted endoscopic group $H={\rm SL}_{3}$ and explicate some of the constructions discussed above. In this case, as we have noted before, the natural map of dual groups:
$\xi_{H}:H^{\vee}={\rm PGL}_{3}({\mathbb{C}})\longrightarrow{\rm PGSO}_{8}^{\vee}={\rm Spin}_{8}({\mathbb{C}})$
is the unique lift of the adjoint representation
${\rm Ad}:{\rm PGL}_{3}({\mathbb{C}})\longrightarrow{\rm SO}_{8}({\mathbb{C}})$
of ${\rm PGL}_{3}({\mathbb{C}})$ on its Lie algebra, from ${\rm SO}_{8}({\mathbb{C}})$ to ${\rm Spin}_{8}({\mathbb{C}})$.
Let us explicate explicate the map
$\xi:T_{H}^{\vee}\longrightarrow T^{\vee}$
and its dual
$\eta:T\longrightarrow T_{H}\subset{\rm SL}_{3}.$
We shall regard $T_{H}^{\vee}$ as the diagonal torus in ${\rm PGL}_{3}({\mathbb{C}})$ and recall from the discussion in §3.1 that the torus $T^{\vee}$ is given by
$T^{\vee}=\{\alpha_{0}(u)\cdot\alpha_{1}(t_{1})\cdot\alpha_{2}(t_{2})\cdot\alpha_{3}(t_{3}):u,t_{i}\in\mathbb{G}_{m}\}.$
Then the map $\xi$ is given on the diagonal maximal torus by:
$\xi:\left(\begin{array}[]{ccc}a&&\\ &b&\\ &&c\end{array}\right)\mapsto u\left(\frac{ab}{c^{2}},\frac{a}{c}\right)=\alpha_{0}\left(\frac{ab}{c^{2}}\right)\cdot\alpha_{1}\left(\frac{a}{c}\right)\cdot\alpha_{2}\left(\frac{a}{c}\right)\cdot\alpha_{3}\left(\frac{a}{c}\right),$
where the element $s(u,t)\in(T^{\vee})^{\theta}$ was defined in (3.1). On the other hand , we may realize $T_{H}$ as the diagonal torus in $H={\rm SL}_{3}$, and identify the torus $T$ with $\mathbb{G}_{m}^{4}$ via:
$T=\{\omega_{0}^{\vee}(u)\cdot\omega_{1}^{\vee}(t_{1})\cdot\omega_{2}^{\vee}(t_{2})\cdot\omega_{3}^{\vee}(t_{3}):u,t_{i}\in\mathbb{G}_{m}\},$
where $\{\omega_{0}^{\vee},\omega_{1}^{\vee},\omega_{2}^{\vee},\omega_{3}^{\vee}\}$ is the basis of $X_{*}(T)$ formed by fundamental coweights, so that $\langle\alpha_{i},\omega_{j}^{\vee}\rangle=\delta_{ij}$. Then the dual map $\eta:T\longrightarrow T_{H}$ is given by
$\eta:\omega_{0}^{\vee}(u)\cdot\omega_{1}^{\vee}(t_{1})\cdot\omega_{2}^{\vee}(t_{2})\cdot\omega_{3}^{\vee}(t_{3})\mapsto\left(\begin{array}[]{ccc}ut_{1}t_{2}t_{3}&&\\ &u&\\ &&u^{-2}t_{1}^{-1}t_{2}^{-1}t_{3}^{-1}\end{array}\right).$
The kernel of $\eta$ is thus
${\rm Ker}(\eta)=\{\omega_{1}^{\vee}(t_{1})\cdot\omega_{2}^{\vee}(t_{2})\cdot\omega_{3}^{\vee}(t_{3}):t_{1}t_{2}t_{3}=1\}={}^{1-\theta}T\simeq\mathbb{G}_{m}^{2}.$
We are not sure if these explicit formulas are especially attractive, but perhaps they provide a certain level of psychological comfort and may certainly be useful for explicit computations.
5.5. Fibers of unramified transfer
The unramified spectral transfer defined by the twisted endoscopic character identity established in Proposition 5.4 realizes the functorial lifting of unramified L-packets associated to $\xi$. Since unramified L-packets are parametrized by semsimple conjugacy classes in the dual group, this unramified spectral transfer is thus specififed by the induced map
$\xi:T_{H}^{\vee}({\mathbb{C}})/W(T_{H}^{\vee})\longrightarrow T^{\vee}({\mathbb{C}})/W(T^{\vee})$
on Weyl-orbits on maximal (dual) tori.
The following proposition determines the nonempty fibers of the map $\xi$.
Proposition 5.8.
Suppose that $[s]$ and $s^{\prime}]$ are two semisimple conjugacy classes of ${\rm PGL}_{3}({\mathbb{C}})$ such that $\xi(s)=\xi(s^{\prime})$. Then one has $s^{\prime}]=[s]$ or $s^{\prime}]=[s]^{-1}$. In particular,
$\Theta_{s}\circ{\rm trans}_{{\rm SL}_{3}}=\Theta_{s^{\prime}}\circ{\rm trans}_{{\rm SL}_{3}}\Longleftrightarrow\text{ $[s^{\prime}]=[s]$ or $[s^{\prime}]=[s]^{-1}$.}$
Proof.
If $\xi(s)=\xi(s^{\prime})$, then under the adjoint representation of ${\rm PGL}_{3}({\mathbb{C}})$, the elements ${\rm Ad}(s)$ and ${\rm Ad}(s^{\prime})$ of ${\rm SO}_{8}({\mathbb{C}})$ have the same set of eigenvalues. The eigenvalues of ${\rm Ad}(s)$ are precisely the value $\gamma(s)$ of the 6 roots $\gamma$ of ${\rm PGL}_{3}({\mathbb{C}})$ on $s$, together with the eigenvalue 1 (with multiplicity 2). Now if we group the 6 roots in 3 pairs (by grouping $\pm\gamma$ together), then the Weyl group $S_{3}$ of ${\rm PGL}_{3}({\mathbb{C}})$ acts on the 3 pairs of roots by permutation. By applying an appropriate Weyl element, we may thus assume that for any pair of roots $\pm\gamma$,
$\{\gamma(s),\gamma(s)^{-1}\}=\{\gamma(s^{\prime}),\gamma(s^{\prime})^{-1}\}.$
We need to show that this forces $s^{\prime}]=[s]$ or $s^{-1}]$.
Suppose that $\alpha$ and $\beta$ are the two simple roots, with $\alpha+\beta$ the highest root. We consider two cases:
  • (a)
    If
    $\alpha(s^{\prime})=\alpha(s)^{\epsilon}\quad\text{and}\quad\beta(s^{\prime})=\beta(s)^{\epsilon}$
    for a fixed $\epsilon=\pm 1$, then we conclude that $s^{\prime}=s^{\epsilon}$, as desired.
  • (b)
    If
    $\alpha(s^{\prime})=\alpha(s)\quad\text{and}\quad\beta(s^{\prime})=\beta(s)^{-1},$
    then we have
    $(\alpha+\beta)(s^{\prime})=\alpha(s)\cdot\beta(s)^{-1}.$
    But we also know that
    $(\alpha+\beta)(s^{\prime})=(\alpha+\beta)(s)\quad\text{or}\quad(\alpha+\beta)(s^{\prime})^{-1}.$
    This implies, for the two respective cases, that
    $\beta(s)=\epsilon\quad\text{or}\quad\alpha(s)=\epsilon$
    for some $\epsilon=\pm 1$. But then, in the respective case, we have
    $\beta(s^{\prime})=\beta(s)^{-1}=\beta(s)\quad\text{or}\quad\alpha(s^{\prime})=\alpha(s)=\alpha(s)^{-1}.$
    In any case, we are back to the scenario (a).
This proves the proposition. ∎
We can provide another perspective on the above proposition. As we noted in Proposition 4.6(iv), the map $\xi:{\rm PGL}_{3}({\mathbb{C}})\rightarrow{\rm Spin}_{8}({\mathbb{C}})$ gives rise to a homomorphism of Hecke algebras
$\xi^{*}:\mathcal{H}_{{\rm PGSO}_{8}}\longrightarrow\mathcal{H}_{{\rm SL}_{3}}.$
Indeed, via the Satake isomorphism, we have
$\mathcal{H}_{{\rm PGSO}_{8}}\simeq{\rm R}[{\rm Spin}_{8}({\mathbb{C}})]\quad\text{and}\quad\mathcal{H}_{{\rm SL}_{3}}\simeq{\rm R}[{\rm PGL}_{3}({\mathbb{C}})]$
where ${\rm R}[G]$ denotes the representation ring of finite-dimensional algebraic representations of $G$. The map $\xi^{*}$ is then given by the restriction of representations via $\xi$. Now ${\rm R}[{\rm Spin}_{8}({\mathbb{C}})]$ is generated by the fundamental representations, namely its three 8-dimensional representations $\{\rho_{1},\rho_{2},\rho_{3}\}$ and the adjoint representation $\mathfrak{so}_{8}$, whereas ${\rm R}[{\rm PGL}_{3}({\mathbb{C}})]$ is generated by the adjoint representation $\mathfrak{sl}_{3}$ as well as ${\rm Sym}^{3}({\rm std})$ and ${\rm Sym}^{3}({\rm std}^{\vee})$, where $\rm std$} denotes the standard 3-dimensional representation of ${\rm SL}_{3}({\mathbb{C}})$. Under restriction via $\xi$, one has
$\rho_{i}\mapsto\mathfrak{sl}_{3}\quad\text{and}\quad\mathfrak{so}_{8}\mapsto\mathfrak{sl}_{3}\oplus{\rm Sym}^{3}({\rm std})\oplus{\rm Sym}^{3}({\rm std}^{\vee}).$
From this , we see that the restriction map $\xi^{*}$ is not surjective. Indeed, its image is contained in the subring ${\rm R}[{\rm PGL}_{3}({\mathbb{C}})]^{+}$ of self-dual representations of ${\rm PGL}_{3}({\mathbb{C}})$. This is the subring fixed by the nontrivial outer automorphism of ${\rm PGL}_{3}({\mathbb{C}})$. This explains why both $\chi\circ\xi^{*}=\chi^{-1}\circ\xi^{*}$ for any character $\chi:\mathcal{H}_{{\rm SL}_{3}}\rightarrow{\mathbb{C}}$.
6. Cuspidal Spectrum of ${\rm SL}_{n}$
Since we are interested in showing the twisted endoscopic transfer from ${\rm SL}_{3}$ to ${\rm PGSO}_{8}$, it will be pertinent to recall some facts about the cuspidal automorphic spectrum of ${\rm SL}_{3}$. We shall consider the case of ${\rm SL}_{n}$ below, though we shall only be interested in the case $n=3$ in this paper. Our discussion in this section is informed by the papers of Labesse-Langlands [LL], Blasius [B], Lapid [L1] and Hiraga-Saito [HiS1, HiS2].
6.1. Local L-packets
Let $F$ be a local field. The L-packets of ${\rm SL}_{n}(F)$ are defined as the set of irreducible constituents of the restriction of an irreducible representation $\pi$ of ${\rm GL}_{n}(F)$. This restriction is known to be semisimple and multiplicity-free [GK] and we denote the corresponding L-packet of ${\rm SL}_{n}(F)$ by $\mathcal{L}[\pi]\subset{\rm Irr}({\rm SL}_{n}(F))$. Moreover, for irreducible representations $\pi$ and $\pi^{\prime}$ of ${\rm GL}_{n}(F)$, one has
$\mathcal{L}[\pi^{\prime}]\cap\mathcal{L}[\pi]\neq\emptyset$
if and only if
$\pi^{\prime}\simeq\pi\otimes(\chi\circ\det)$ for some character $\chi$ of $F^{\times}$,
in which case one has $\mathcal{L}[\pi^{\prime}]=\mathcal{L}[\pi]\subset{\rm Irr}({\rm SL}_{n}(F))$. We shall also write $\mathcal{L}[\pi]$ for the multiplicity-free semisimple representation given by the direct sum of the members of $\mathcal{L}[\pi]$.
In particular, any unramified L-packet of ${\rm SL}_{n}(F)$ is of the form $\mathcal{L}[\pi]$ for $\pi$ an unramified representation of ${\rm GL}_{n}(F)$. If the Satake parameter of $\pi$ is the semisimple conjugacy class $c(\pi)\in{\rm GL}_{n}({\mathbb{C}})$, then that for $\mathcal{L}[\pi]$ is the image of $c(\pi)$ under the natural projection map
$p:{\rm GL}_{n}({\mathbb{C}})\longrightarrow{\rm PGL}_{n}({\mathbb{C}})={\rm SL}_{n}^{\vee}.$
More generally, if $\phi_{\pi}$ is the L-parameter of $\pi\in{\rm Irr}({\rm GL}_{n}(F))$, then the L-parameter of $\mathcal{L}[\pi]$ under the local Langlands correspondence for ${\rm SL}_{n}(F)$ is given by $p\circ\phi_{\pi}$.
6.2. Endoscopy
The internal structure of a local L-packet $\mathcal{L}[\pi]$ can be probed via the theory of endoscopy. For this purpose, let us explicate the endoscopic data for ${\rm SL}_{3}$, since this is the main case of interest for us.
Lemma 6.1.
Up to isomorphism, the local endoscopic data for ${\rm SL}_{3}$ are given by pairs $(E/F,s)$ where
  • $E/F$ is a cyclic (Galois) cubic field extension;
  • $s\in{\rm PGL}_{3}({\mathbb{C}})={\rm SL}_{3}^{\vee}$ is one of the following two elements
    (6.2) $\left(\begin{array}[]{ccc}1&&\\ &\zeta&\\ &&\zeta^{2}\end{array}\right)\quad\text{or}\quad\left(\begin{array}[]{ccc}1&&\\ &\zeta^{2}&\\ &&\zeta\end{array}\right).$
For a given $(E/F,s)$, the associated endoscopic group is the norm-one torus
$T_{E}:={\rm Ker}(N_{E/k}:{\rm Res}_{E/k}(\mathbb{G}_{m})\longrightarrow\mathbb{G}_{m})$
and the associated embedding
$\xi_{E/F,s}:{{}^{L}}T_{E}\simeq T_{E}^{\vee}\rtimes{\rm Gal}(E/F)\longrightarrow{\rm PGL}_{3}({\mathbb{C}})$
can be described as follows.
In particular, for each cyclic cubic extension $E/F$, the norm-one torus $T_{E}$ occurs as an endoscopic group two times; in other words, there are two transfer maps
${\rm trans}_{T_{E},s}:\mathcal{I}({\rm SL}_{3})\longrightarrow\mathcal{SI}(T_{E})=\mathcal{I}(T_{E})$
corresponding to the two choices for $s$.
6.3. Endoscopic local L-packets
For general $n$, it has been shown by Henniart-Herb [HH], Hiraga-Saito [HiS1, HiS2] and Hiraga-Ichino [HI] that the L-packets of ${\rm SL}_{n}(F)$, defined in the above manner, satisfies the expected endoscopic character identities, as dictated by the theory of endoscopy. Hence, the span of the Harish-Chandra characters of the members of $\mathcal{L}[\pi]$ contains a 1-dimensional subspace of stable distribution, given by the sum of all the characters of all members of $\mathcal{L}[\pi]$ and hence is simply given by the Harish-Chandra character $\Theta_{\pi}$ of $\pi$. We shall refer to $\Theta_{\pi}$ as the stable trace of the L-packet $\mathcal{L}[\pi]$ and denote it by $S\Theta_{[\pi]}$.
Let us illustrate with the case of endoscopic local L-packets of ${\rm SL}_{3}$. These endoscopic local L-packets are those whose L-parameter factors through some $\xi_{E/F,s}$. They are precisely the $\mathcal{L}[\pi]$ such that $\pi\otimes\omega_{E/F}\simeq\pi$ for some cyclic cubic extension $E/F$ with associated cubic character $\omega_{E/F}^{\pm 1}$. In particular, the L-parameter of $\pi$ is of the form ${\rm Ind}_{W_{E}}^{W_{F}}\chi$ for some character $\chi$ of $E^{\times}$, in which case we will write $\pi=\pi_{\chi}$.
Let us assume that $\chi$ is not fixed by ${\rm Gal}(E/F)$, so that $\pi_{\chi}$ is supercuspidal. For a generic choice of $\chi$, $\mathcal{L}[\pi_{\chi}]$ is the sum of 3 irreducible representations, say $\pi_{1}$, $\pi_{2}$ and $\pi_{3}$. Then the local endoscopic character identities expresses ${\rm trans}_{T_{E},s}(\chi)$ (for the two choices of $s$) as a linear combination of the Harish-Chandra characters of the $\pi_{i}$’s. More precisely, one has
$\begin{cases}{\rm trans}_{T_{E},s}(\chi)=\Theta_{\pi_{1}}+\zeta\cdot\Theta_{\pi_{2}}+\zeta^{2}\cdot\Theta_{\pi_{3}};\\ {\rm trans}_{T_{E},s^{-1}}(\chi)=\Theta_{\pi_{1}}+\zeta^{2}\cdot\Theta_{\pi_{2}}+\zeta\cdot\Theta_{\pi_{3}}.\end{cases}$
These two invariant distributions, together with the stable trace $S\Theta_{[\pi_{\chi}]}$, give a basis for the 3-dimensional space of invariant distributions spanned by the characters of the $\pi_{i}$’s.
6.4. Cuspidal spectrum
Now let $k$ be a number field. If $\pi\simeq\bigotimes^{\prime}_{v}\pi_{v}\subset\mathcal{A}_{cusp}({\rm GL}_{n})$ is a cuspidal automorphic representation of ${\rm GL}_{n}$, the abstract restriction of $\pi$ to ${\rm SL}_{n}({\mathbb{A}})$ gives rise to a multiplicity-free semisimple representation of ${\rm SL}_{n}({\mathbb{A}})$, all of whose irreducible summands are nearly equivalent with each other. This is a global L-packet of ${\rm SL}_{n}$ and we shall denote this by $\mathcal{L}[\pi]$, noting that
$\mathcal{L}[\pi]\simeq\bigotimes^{\prime}_{v}\mathcal{L}[\pi_{v}].$
On the other hand, one has the restriction of automorphic forms
${\rm rest}:\mathcal{A}_{cusp}({\rm GL}_{n})\longrightarrow\mathcal{A}_{cusp}({\rm SL}_{n}).$
Then the image $L[\pi]:={\rm rest}(\pi)$ is a submodule of both $\mathcal{A}_{cusp}({\rm SL}_{n})$ and $\mathcal{L}[\pi]$ as an ${\rm SL}_{n}({\mathbb{A}})$-module.
As the restriction map $\rm rest$} is surjective, one can hope to understand the cuspidal spectrum of ${\rm SL}_{n}$ in terms of that of ${\rm GL}_{n}$. For this purpose, following Lapid [L1], we consider the following three equivalence relations on cuspidal representations of ${\rm GL}_{n}$:
Definition 6.3.
Let $\pi$ and $\pi^{\prime}$ be cuspidal representations of ${\rm GL}_{n}$.
(a) Write $\pi{\thicksim}_{s}\pi^{\prime}$ if and only if
$\pi^{\prime}\simeq\pi\otimes(\chi\circ\det)$ for some character $\chi$ of $F^{\times}\backslash{\mathbb{A}}^{\times}$,
in which case one has
$\pi^{\prime}=\pi\otimes(\chi\circ\det)\subset\mathcal{A}_{cusp}({\rm GL}_{n}).$
We shall denote the equivalence class of $\pi$ under this equivalence relation by $\pi]_{s}$.
(b) Write $\pi{\thicksim}_{ew}\pi^{\prime}$ if and only if
$\pi^{\prime}\simeq\pi\otimes(\chi\circ\det)$ for some character $\chi$ of $\mathbb{A}$$^{\times}$.
We shall denote the equivalence class of $\pi$ under this equivalence relation by $\pi]_{ew}$.
(c) Write $\pi{\thicksim}_{w}\pi^{\prime}$ if and only if
$\pi^{\prime}$ is nearly equivalent to $\pi\otimes(\chi\circ\det)$ for some character $\chi$ of $\mathbb{A}$$^{\times}$.
In other words,
$\pi^{\prime}_{v}\simeq\pi_{v}\otimes(\chi_{v}\circ\det)$ for some character $\chi_{v}$ outside a finite set $S$ of places of $k$.
We shall denote the equivalence class of $\pi$ under this equivalence relation by $\pi]_{w}$.
Clearly, the equivalence relation gets weaker as one goes down the list. When $n=2$, it is a result of Ramakrishnan [Ra1] that the three equivalence relations above are in fact the same. For $n\geq 3$, Blasius [B] has shown that they are distinct from each other. The relevance of the three equivalence relations to the cuspidal spectrum of ${\rm SL}_{n}$ is addressed by the following proposition [L1, HiS1, HiS2]:
Proposition 6.4.
Suppose that $\pi$ and $\pi^{\prime}$ are cuspidal representations of ${\rm GL}_{n}$. Then we have:
(i) $\pi{\thicksim}_{s}\pi^{\prime}$ if and only if
$0\neq L[\pi^{\prime}]\cap L[\pi]\subset\mathcal{A}_{cusp}({\rm SL}_{n}),$
in which case one has $L[\pi^{\prime}]=L[\pi]$.
(ii) $\pi{\thicksim}_{ew}\pi^{\prime}$ if and only if
$L[\pi^{\prime}]\simeq L[\pi]$ as abstract ${\rm SL}_{n}({\mathbb{A}})$-modules.
(iii) $\pi{\thicksim}_{w}\pi^{\prime}$ if and only if
$L[\pi]$ and $L[\pi^{\prime}]$ have nearly equivalent irreducible summands.
Observe that when $\pi$ and $\pi^{\prime}$ satisfies (iii) but not (ii), then $\mathcal{L}[\pi]$ and $\mathcal{L}[\pi^{\prime}]$ have no isomorphic irreducible summands, because their local components at some places belong to different local L-packets, and hence are inequivalent. In other words, the so-called strong multiplicity one property (or more appropriately, rigidity) does not hold for ${\rm SL}_{n}$, at the level of L-packets.
As a consequence of the above proposition, one obtains the following decomposition of the cuspidal spectrum of ${\rm SL}_{n}$:
$\mathcal{A}_{cusp}({\rm SL}_{n})=\bigoplus_{[\pi]_{s}}L[\pi]=\bigoplus_{[\pi]_{ew}}\left(\bigoplus_{[\pi^{\prime}]_{s}{\thicksim}_{ew}[\pi]_{s}}L[\pi^{\prime}]\right)\simeq\bigoplus_{[\pi]_{ew}}m_{\mathcal{L}[\pi]}L[\pi].$
Here, the first sum runs over the ${\thicksim}_{s}$-equivalence classes $\pi]_{s}$ of cuspidal representations of ${\rm GL}_{n}$, whereas the subsequent sums run over the ${\thicksim}_{ew}$-equivalence classes $\pi]_{ew}$. Moreover, $m_{\mathcal{L}[\pi]}$ is called the multiplicity of the L-packet $\mathcal{L}[\pi]$ and is equal to the number of ${\thicksim}_{s}$-equivalence classes in the ${\thicksim}_{ew}$-equivalence class of $\pi$. In particular, $m_{\mathcal{L}[\pi]}$ gives the multiplicity of the irreducible summands of $L[\pi]$ in the cuspidal spectrum of ${\rm SL}_{n}$.
We may also write
$\mathcal{A}_{cusp}({\rm SL}_{n})=\bigoplus_{[\pi]_{w}}\left(\bigoplus_{[\pi^{\prime}]_{s}{\thicksim}_{w}[\pi]_{s}}L[\pi^{\prime}]\right).$
where the outer sum runs over the ${\thicksim}_{w}$-equivalence classes of cuspidal representations of ${\rm GL}_{n}$ and the inner sum runs over the ${\thicksim}_{s}$-equivalence classes in the ${\thicksim}_{w}$-equivalence class of $\pi$. This is a decomposition of the cuspidal spectrum of ${\rm SL}_{n}$ into near equivalence classes of representations of ${\rm SL}_{n}({\mathbb{A}})$. Following Lapid, we define the global multiplicity of $\mathcal{L}[\pi]$ by:
(6.5) $\mathcal{M}_{\mathcal{L}[\pi]}=\sum_{[\pi^{\prime}]_{s}{\thicksim}_{w}[\pi]_{s}}m_{\mathcal{L}[\pi^{\prime}]}.$
6.5. Stable trace formula
Let us now specialize to the case $n=3$. One can enhance the above understanding of $\mathcal{A}_{cusp}({\rm SL}_{3})$ by considering the stable trace formula of ${\rm SL}_{3}$. We give a very brief impressionistic sketch of this.
Analogous to the local case, the elliptic endoscopic data of ${\rm SL}_{3}$ are indexed by pairs $(E/k,s)$ where $E/k$ is a cyclic cubic extension and $s$ is one of the two elements of ${\rm PGL}_{3}({\mathbb{C}})$ given in (6.2). For a given pair $(E/k,s)$, the corresponding elliptic endoscopic group of ${\rm SL}_{3}$ is the anisotropic norm-one torus
$T_{E}:={\rm Ker}(N_{E/k}:{\rm Res}_{E/k}(\mathbb{G}_{m})\longrightarrow\mathbb{G}_{m}).$
Then the stable trace formula of ${\rm SL}_{3}$ takes the form
$I^{{\rm SL}_{3}}_{disc}(f)=S_{disc}^{{\rm SL}_{3}}(f)+\frac{1}{9}\cdot\sum_{(E/k,s)}I_{disc}^{T_{E}}(f_{T_{E},s})$
for test functions $f$ on ${\rm SL}_{3}({\mathbb{A}})$ with transfer $f_{T_{E},s}$ on $T_{E}({\mathbb{A}})$ relative to $(E/k,s)$. Here, $I_{disc}^{{\rm SL}_{3}}$ and $I_{disc}^{T_{E}}$ refers to the discrete part of the invariant trace formula for ${\rm SL}_{3}$ and $T_{E}$ respectively. On the other hand, $S_{disc}^{{\rm SL}_{3}}$ is a stable distribution on ${\rm SL}_{3}({\mathbb{A}})$.
Suppose that $c=\{c_{v}\}_{v\notin S}\subset{\rm PGL}_{3}({\mathbb{C}})$ is a Hecke-Satake family, i.e. a collection of semisimple conjugacy classes in ${\rm PGL}_{3}({\mathbb{C}})$ for all $v$ outside a finite set $S$ of places of $k$. Then $c$ determines a character
$\chi_{c}:\mathcal{H}_{{\rm SL}_{3}}^{S}:=\bigotimes_{v\notin S}\mathcal{H}_{{\rm SL}_{3}(k_{v})}\longrightarrow{\mathbb{C}}$
of the spherical Hecke algebras outside $S$ and hence a near equivalence class of representations of ${\rm SL}_{3}({\mathbb{A}})$. Considering the $c$-isotypic part of the stable trace formula, i.e. where the Hecke algebra $\mathcal{H}_{{\rm SL}_{3}}^{S}$ acts by the character $\chi_{c}$, we obtain the more refined identity:
$I_{disc,c}^{{\rm SL}_{3}}(f_{S})=S_{disc,c}^{{\rm SL}_{3}}(f_{S})+\frac{1}{9}\cdot\sum_{(E/k,s)}I_{disc,c}^{T_{E}}({\rm trans}_{E/k,s}(f_{S})),\quad\text{for $f_{S}\in\bigotimes_{v\in S}C^{\infty}_{c}({\rm SL}_{3}(k_{v}))$,}$
where $I_{disc,c}^{{\rm SL}_{3}}$ and $S_{disc,c}^{{\rm SL}_{3}}$ denote the $c$-isotypic part of the relevant distributions, i.e. where the Hecke algebra $\mathcal{H}_{{\rm SL}_{3}}^{S}$ acts by the character $\chi_{c}$.
6.6. Endoscopic L-packets
A cuspidal global L-packet $\mathcal{L}[\pi]$ of ${\rm SL}_{3}$ is endoscopic if its Hecke-Satake family $c(\pi)$ satisfies
$I_{disc,c}^{T_{E}}\circ{\rm trans}_{E/k,s}\neq 0\quad\text{ for some $(E/k,s)$.}$
Such an L-packet is thus the endoscopic transfer of a nontrivial automorphic character of $T_{E}$.
To describe these endoscopic L-packets more explicitly, let us start with a nontrivial automorphic character $\chi$ of ${\mathbb{A}}_{E}^{\times}$ and consider the cuspidal representation $\pi_{\chi}$ of ${\rm GL}_{3}$ obtained by automorphic induction from $\chi$ (see §8.2 below for a more detailed discussion of automorphic induction). The cuspidal representations of ${\rm GL}_{3}$ which are obtained by automorphic induction from a cyclic cubic extension are characterized as those which are invariant under twisting by an automorphic (nontrivial) cubic character (corresponding to the cyclic cubic extension in question). Then the endoscopic L-packet of ${\rm SL}_{3}$ determined by $\chi|_{T_{E}}$ is simply $\mathcal{L}[\pi_{\chi}]$, so that its contribution to the cuspidal spectrum of ${\rm SL}_{3}$ is the submodule $L[\pi_{\chi}]$.
Observe that the property of a global L-packet $\mathcal{L}[\pi]$ being endoscopic depends only on the ${\thicksim}_{w}$-equivalence class of the cuspidal representation $\pi$ of ${\rm GL}_{3}$. Indeed, if $\pi^{\prime}{\thicksim}_{w}\pi$, then there is an abstract character $\mu$ of $\mathbb{A}$$^{\times}$ such that $\pi^{\prime}$ and $\pi\otimes\mu$ are nearly equivalent. Hence, if $\pi$ is invariant under twisting by a nontrivial cubic automorphic character $\nu$, then $\pi^{\prime}\otimes\nu$ and $\pi^{\prime}$ are nearly equivalent. By the strong multiplicity or rigidity property, we must have $\pi^{\prime}\otimes\nu\simeq\pi^{\prime}$, so that $\pi^{\prime}$ is obtained as an automorphic induction from the same cubic field extension as $\pi$.
6.7. Non-endoscopic L-packets
On the other hand, suppose that $\mathcal{L}[\pi]$ is not endoscopic, i.e.
$I_{disc,c(\pi)}^{T_{E}}\circ{\rm trans}_{E/k,s}\equiv 0\quad\text{for any $(E/k,s)$.}$
In this case, one has
$S_{disc,c(\pi)}^{{\rm SL}_{3}}=I_{disc,c(\pi)}^{{\rm SL}_{3}}=\bigoplus_{[\pi^{\prime}]_{s}{\thicksim}_{w}[\pi]_{s}}L[\pi^{\prime}].$
Since this distribution is stable, we deduce that
$L[\pi^{\prime}]\simeq\mathcal{L}[\pi^{\prime}]\quad\text{for each $[\pi^{\prime}]_{s}{\thicksim}_{w}[\pi]_{s}$,}$
so that each member of the global L-packet $\mathcal{L}[\pi^{\prime}]$ is cuspidal automoprhic. Hence we deduce that
(6.6) $S_{disc,c(\pi)}^{{\rm SL}_{3}}=\bigoplus_{[\pi^{\prime}]_{s}{\thicksim}_{w}[\pi]_{s}}S\Theta_{[\pi^{\prime}]},$
where
$S\Theta_{[\pi^{\prime}]}=\bigotimes_{v}S\Theta_{[\pi^{\prime}_{v}]}=\Theta_{\pi^{\prime}}|_{{\rm SL}_{3}({\mathbb{A}})}\quad\text{ if $\pi^{\prime}\simeq\otimes^{\prime}_{v}\pi^{\prime}_{v}$.}$
6.8. Stable multiplicity formula
By the stable trace formula recounted above, the local endoscopic character identities mentioned in §6.3 and the description of the cuspidal spectrum of ${\rm SL}_{3}$ given in §6.4, one can obtain the same formula (6.6) for the endoscopic L-packets. As a consequence, one has an explicit spectral decomposition of the cuspidal part of $S_{disc}^{{\rm SL}_{3}}$, namely
(6.7) $S_{cusp}^{{\rm SL}_{3}}=\bigoplus_{[\pi]_{s}}S\Theta_{[\pi]},$
where the sum runs over the ${\thicksim}_{s}$-equivalence classes $\pi]_{s}$ of cuspidal automorphic representations $\pi$ of ${\rm GL}_{3}$. This is sometimes called the stable multiplicity formula for ${\rm SL}_{3}$.
7. Adjoint lifting of ${\rm GL}_{3}$
In this section, we exploit the stable twisted trace formula for the twisted space $({\rm PGSO}_{8},\theta)$ to show the adjoint lifting from ${\rm GL}_{3}$ to ${\rm GL}_{8}$. Recall from the introduction that the adjoint representation ${\rm Ad}:{\rm GL}_{3}({\mathbb{C}})\longrightarrow{\rm GL}_{8}({\mathbb{C}})$ factors as:
$\begin{CD}{\rm GL}_{3}({\mathbb{C}})@>{}>{}>{\rm PGL}_{3}({\mathbb{C}})@>{\tilde{{\rm Ad}}}>{}>{\rm Spin}_{8}({\mathbb{C}})@>{}>{}>{\rm SO}_{8}({\mathbb{C}})@>{\rm std}>{}>{\rm GL}_{8}({\mathbb{C}}),\end{CD}$
The functoriality relative to all other arrows being known, our purpose in this section is to establish the functorial lifting with respect to $\tilde{\rm Ad}$.
7.1. Global K-forms of ${\rm PGSO}_{8}$
As we noted in §4.1, a slight complication with the theory of stable (twisted) trace formula is the need to take into account of K-forms [MW1, §I.1.11], which are certain inner forms of the twisted space in question. In the context of $({\rm PGSO}_{8},\theta)$, we have seen there that the K-forms correspond to isomorphism classes of octonion algebras over local or number fields. Thus, over a non-Archimedean local field or $\mathbb{C}$}, $({\rm PGSO}_{8},\theta)$ has no additional $K$-forms, since there is a unique octonion algebra up to isomorphism. On the other hand, over $\mathbb{R}$}, there are two octonions algebras: the split one and a division algebra. As such, there are two K-forms of $({\rm PGSO}_{8},\theta)$ over $\mathbb{R}$}. For a number field $k$, with $r$ real places, the number of such octonion algebras is $2^{r}$, so that there are $2^{r}$ global K-forms over $k$.
7.2. Stable twisted trace formula.
The stable twisted trace formula for $({\rm PGSO}_{8},\theta)$ takes the form [MW2, Thm. X.8.1, Pg 1241]
(7.1) $\sum_{\mathbb{O}}I^{\mathbb{O},\theta}_{disc}(f_{\mathbb{O}})=S^{{\rm G}_{2}}_{disc}(f_{{\rm G}_{2}})+\frac{1}{2}S_{disc}^{{\rm SO}_{4}}(f_{{\rm SO}_{4}})+S_{disc}^{{\rm SL}_{3}}(f_{{\rm SL}_{3}}),$
where $\{f_{\mathbb{O}}\in C^{\infty}_{c}({\rm PGSO}^{\mathbb{O}}_{8}({\mathbb{A}}))\}$ is a family of test functions on the K-forms and $f_{{\rm G}_{2}}$, $f_{{\rm SO}_{4}}$ and $f_{{\rm SL}_{3}}$ are its twisted endoscopic transfer to the twisted endoscopic groups ${\rm G}_{2}$, ${\rm SO}_{4}$ and ${\rm SL}_{3}$ respectively. For each twisted endoscopic group $H$, the stable distribution $S^{H}_{disc}$ is the discrete part of the stable trace formula for $H$.
Let us explicate the distributions $I^{\mathbb{O},\theta}_{disc}$, following [LW, §14.3] and [MW2, §X.5.1]. Writing $G={\rm PGSO}^{\mathbb{O}}_{8}$, one has the expression (see [LW, Thm. 14.3.1 and Prop. 14.3.2] and [A, Pg. 125-128]).
$I^{\mathbb{O},\theta}_{disc}(f)=\sum_{M}\frac{1}{|W(G,M)|}I^{\mathbb{O},\theta}_{disc,M}(f)\quad\text{ for $f\in C^{\infty}_{c}(G({\mathbb{A}}))$, }$
where the sum runs over the conjugacy classes of Levi subgroups $M$ of $G$ and [MW2, §X.5.1]
$I^{\mathbb{O},\theta}_{disc,M}(f)=\sum_{s\in W(G\cdot\theta,M)_{reg}}|\det(1-s|\mathfrak{a}_{M})|^{-1}\cdot{\rm Tr}(M_{P|s(P)}(0)\cdot\rho_{P,s,disc}(f)).$
Here:
    $P=M\cdot U$ is a standard parabolic subgroup with Levi subgroup $M$ and unipotent radical $U$;
    $W(G\cdot\theta,M)$ is the Weyl set of elements of $G(F)\cdot\theta$ which normalizes $M$ (taken modulo $M(F)$), and $W(G\cdot\theta,M)_{reg}$ is the subset of regular elements, i.e. those $s$ such that $\det(1-s|\mathfrak{a}_{M})\neq 0$.
  • we regard elements $s\in G(F)\cdot\theta$ as outer automorphisms of $G$;
    $\rho_{P,s,disc}$ refers to the representation of the twisted space $G({\mathbb{A}})\cdot\theta$ on
[…]
    with the action of $\gamma\cdot\theta\in G({\mathbb{A}})\cdot\theta$ given as the composite:
[…]
    where
    $(r_{s}(\gamma\cdot\theta)\phi)(g)=\phi(s^{-1}(g\gamma)).$
    and $M_{P,s\cdot P}$ is a standard intertwining operator.
  • for $f\in C^{\infty}_{c}(G({\mathbb{A}}))$, one has
    $\rho_{P,s,disc}(f)=\int_{G({\mathbb{A}})}f(\gamma)\cdot\rho_{P,s,disc}(\gamma\cdot\theta)\,d\gamma.$
In particular, when $M=G$, one has $s=\theta$ and
$\rho_{G,\theta,disc}(\gamma\cdot\theta)\phi(g)=\phi(\theta^{-1}(g\gamma)),\quad\text{ for $\phi\in\mathcal{A}_{disc}(G)$.}$
Thus,
$I^{\mathbb{O},\theta}_{disc,G}(f)={\rm Tr}(f\cdot\theta|\mathcal{A}_{disc}(G)),$
where the action of $f\in C^{\infty}_{c}(G({\mathbb{A}}))$ on $\mathcal{A}_{disc}(G))$ is the usual one and that of $\theta$ is given by
$\theta(\phi)(g)=\phi(\theta^{-1}(g)).$
On the other hand, for $M\neq G$, the representations of $G({\mathbb{A}})$ which intervene in $I^{\mathbb{O},\theta}_{disc,M}$ are the $\theta$-stable representations of the form
${\rm Ind}_{P({\mathbb{A}})}^{G({\mathbb{A}})}\sigma,$
where $\sigma\subset\mathcal{A}_{disc,\chi}(M)$ is a $s$-stable submodule with unitary central character $\chi$, for some $s\in W(G\cdot\theta,M)_{reg}$.
7.3. Adjoint lifting.
Using the stable twisted trace formula recalled above, we can now show the following theorem:
Theorem 7.2.
Let $\pi$ be a cuspidal representation of ${\rm GL}_{3}$ such that $\pi_{v_{0}}$ is a discrete series representation for some place $v_{0}$. Then the (weak) adjoint lifting of $\pi$ to ${\rm GL}_{8}$ exists.
Proof.
To use the stable twisted trace formula (7.1), we need to specify the test function $f_{\mathbb{O}}=\prod_{v}f_{\mathbb{O},v}$ used there, for each octonion $k$-algebra $\mathbb{O}$. We do this as follows:
  • At a real place $v$, we let the test functions $f_{\mathbb{O},v}$ be arbitrary (so that they may depend on $\mathbb{O}$).
  • If $v$ is a non-Archimedean distinct from $v_{0}$ or a complex place of $k$, we will choose $f_{\mathbb{O},v}=f_{v}$ to be independent of $\mathbb{O}$ but otherwise arbitrary.
  • At the place $v_{0}$, we take a a pseudo-coefficient $f_{\pi_{v_{0}}}$ of $\pi_{v_{0}}$ and consider its stable orbital integral which is a nonzero element of $\mathcal{SI}_{cusp}({\rm SL}_{3})$. By Proposition 4.6(i), one then pick $f_{v_{0}}\in\mathcal{I}^{\theta}_{cusp}({\rm PGSO}_{8})$ such that
    (7.3) ${\rm trans}([f_{v_{0}}])=(0,0,[f_{\pi_{v_{0}}}]).$
    We then take $f_{\mathbb{O},v_{0}}=f_{v_{0}}$ for any $\mathbb{O}$.
With such a family of test functions $f_{\mathbb{O}}\}$, (7.1) simplifies to:
[…]
where
$f_{{\rm SL}_{3}}=\sum_{\mathbb{O}}{\rm trans}^{\mathbb{O}}_{{\rm SL}_{3}}(f_{\mathbb{O}})=\left(\prod_{\text{real $v$}}\left(\sum_{\mathbb{O}_{v}}{\rm trans}^{\mathbb{O}_{v}}_{{\rm SL}_{3}}(f_{\mathbb{O},v})\right)\right)\cdot\prod_{\text{other $v$}}{\rm trans}_{{\rm SL}_{3}}(f_{v}),$
with the sum over $\mathbb{O}_{v}$ running over the two octonion algebras at a real place $v$.
Now we may use the action of the spherical Hecke algebras of $G={\rm PGSO}_{8}$ at almost all places of $k$ to isolate a given Hecke-Satake family on the LHS of (7.4). More precisely, with our given cuspidal representation $\pi$ of ${\rm GL}_{3}$ with associated submodule $L[\pi]\subset\mathcal{A}_{cusp}({\rm SL}_{3})$ and Hecke-Satake family
$c([\pi])=\{c([\pi_{v}]):v\notin S\}\subset{\rm PGL}_{3}({\mathbb{C}}),$
let us set
$c:=\{\tilde{\rm Ad}(c([\pi_{v}])):v\notin S\}\subset{\rm Spin}_{8}({\mathbb{C}})$
where $\tilde{\rm Ad}:{\rm PGL}_{3}({\mathbb{C}})\hookrightarrow{\rm Spin}_{8}({\mathbb{C}})$ is the adjoint map and $S$ is a sufficiently large finite set of places, including all Archimedean ones and the place $v_{0}$. At each Archimedean place $v$, we may also consider the infinitesimal character $\lambda([\pi_{v}])$ of $\mathcal{L}[\pi_{v}]$, which may be regarded as a semisimple conjugacy class in the complexified dual Lie algebra $\mathfrak{sl}_{3}$, giving rise to
$\lambda([\pi])=(\lambda([\pi_{v}]))_{v\in S_{\infty}}\in\mathfrak{sl}_{3}(k\otimes_{\mathbb{Q}}{\mathbb{C}}),$
where $S_{\infty}$ denotes the set of Archimedean places of $k$. We set
$\lambda:=(\tilde{\rm Ad}_{*}(\lambda([\pi_{v}])))_{v\in S_{\infty}}\in\mathfrak{so}_{8}(k\otimes_{\mathbb{Q}}{\mathbb{C}}),$
By [MW2, §X.5.9], one has;
(7.5) $\sum_{\mathbb{O}}I_{disc,c,\lambda}^{\mathbb{O},\theta}(f_{\mathbb{O},S})=\sum_{[\pi^{\prime}]_{s}}S^{{\rm SL}_{3}}_{disc,[\pi^{\prime}]_{s}}(f_{{\rm SL}_{3},S})$
where
$f_{\mathbb{O},S}\in\left(\bigotimes_{v\in S}C^{\infty}_{c}({\rm PGSO}_{8}^{\mathbb{O}_{v}})\right)\otimes\left(\bigotimes_{v\notin S}1_{K_{v}}\right)$
with $1_{K_{v}}$ the characteristic function of $K_{v}={\rm PGSO}_{8}(\mathcal{O}_{v})$, and the sum on the RHS ranges over all $\pi^{\prime}]_{s}$ such that
$\tilde{\rm Ad}(c([\pi^{\prime}]))=c\quad\text{and}\quad\tilde{\rm Ad}_{*}(\lambda[\pi^{\prime}])=\lambda.$
We shall denote the set of such $\pi^{\prime}]_{s}$ by $\Xi_{c,\lambda}$.
Now for each $\pi^{\prime}]\in\Xi_{c,\lambda}$, it follows by Proposition 5.8 that, for each $v\notin S$, the local L-packets $\mathcal{L}[\pi^{\prime}_{v}]$ and $\mathcal{L}[\pi_{v}]$ are either equal or contragredient of each other; in particular, $\mathcal{L}[\pi^{\prime}_{v}]$ is a generic L-packet for each $v\notin S$. Moreover, since $f_{v_{0}}$ lies in $\mathcal{I}^{\theta}_{cusp}({\rm PGSO}_{8})$, any $\pi^{\prime}]_{s}\in\Xi_{c,\lambda}$ is necessarily contained in the cuspidal spectrum $\mathcal{A}_{cusp}({\rm SL}_{3})$.
By the stable multiplicity formula (6.7)) for ${\rm SL}_{3}$, (7.5) becomes:
$\sum_{\mathbb{O}}I_{disc,c}^{\mathbb{O},\theta}(f_{\mathbb{O},S})=\sum_{[\pi^{\prime}]_{s}\in\Xi_{c,\lambda}}{\rm Tr}\left(\pi^{\prime}_{S}(f_{{\rm SL}_{3},S})\right).$
Now the RHS is the pullback by ${\rm trans}_{{\rm SL}_{3},S}$ of the character distribution of a stable admissible representation of $\prod_{v\in S}{\rm SL}_{3}(k_{v})$. By Lemma 4.9 (or rather the semi-local version of it), one deduces that the RHS of the above identity defines a nonzero linear functional on $\otimes_{v\in S}C^{\infty}_{c}({\rm PGSO}_{8}^{\mathbb{O}_{v}}(k_{v}))$. This implies that the LHS is a nonzero linear functional and in particular that
$\sum_{\mathbb{O}}I_{disc,c}^{\mathbb{O},\theta}\neq 0.$
This shows that the Hecke-Satake family $c$ can be detected in the spectrum of the $I_{disc}^{\mathbb{O},\theta}$ for some $\mathbb{O}$ and hence can be realized by an automorphic subrepresentation $\Pi_{\mathbb{O}}$ of ${\rm PGSO}_{8}^{\mathbb{O}}$ for some $\mathbb{O}$.
Pulling $\Pi_{\mathbb{O}}$ back to ${\rm SO}_{8}$ via $f_{1}:{\rm SO}_{8}^{\mathbb{O}}\rightarrow{\rm PGSO}_{8}^{\mathbb{O}}$ and transfering of $f_{1}^{*}(\Pi_{\mathbb{O}})$ to ${\rm GL}_{8}$ (appealing to, for example, [CZ] or [CFK] for the case when $\mathbb{O}$ is not split), one obtains the desired weak adjoint lifting of $\pi$. ∎
In fact, it should not be too hard to remove the hypothesis that some local component of $\pi$ is discrete series in the above theorem, using arguments in [A, §3.5]. However, this would involve analyzing contributions from the other elliptic twisted endoscopic groups ${\mathbb{G}}_{2}$ and ${\rm SO}_{4}$. Hence, we do not carry out this more elaborate analysis here. It will be contained in a forthcoming work with S.W. Shin and Y. Takanashi, whose goal is to establish the weak lifting of automorphic representations from the 3 twisted endoscopic group of $({\rm PGSO}_{8},\theta)$ to $\{{\rm PGSO}_{8}^{\mathbb{O}}\}$ and to obtain an endoscopic classification for the discrete automorphic representations of $G_{2}$.
8. Automorphic Induction Case
In this section, we study the case when the cuspidal representation $\pi$ is obtained by automorphic induction from a Hecke character of a cubic field extension $E/k$. In particular, we shall see that the adjoint lifting ${\rm Ad}(\pi)$ always exists (without the local condition in Theorem 7.2. Moreover, we shall be able to determine the possible isobaric decomposition of ${\rm Ad}(\pi)$.
8.1. Cubic fields
Let $E/k$ be a cubic field extension, so that $E/k$ gives rise to a conjugacy class of homomorphsms $\alpha_{E}:{\rm Gal}(\overline{k}/k)\longrightarrow S_{3}$. Composing $\alpha_{E}$ with the sign character of $S_{3}$ gives a quadratic character, which is trivial if and only if $E/k$ is Galois (or normal). The étale quadratic algebra determined by $({\rm sign})\circ\alpha_{E}$ is the discriminant quadratic algebra $K_{E}$ of $E/k$. Composing $\alpha_{E}$ with the two-dimensional irreducible representation of $S_{3}$, one obtains a two-dimensional representation $\rho_{E}$ of ${\rm Gal}(\overline{k}/k)$ which is irreducible if and only if $E/k$ is non-Galois. We note that
${\rm Ind}_{W_{E}}^{W_{k}}1=1\oplus\rho_{E}.$
8.2. Automorphic induction
Given a Hecke character $\chi:E{{}^{\times}}\backslash{\mathbb{A}}_{E}^{\times}\longrightarrow{\mathbb{C}}^{\times}$, one obtains a $1$-dimensional representation (still denoted by $\chi$)
$\chi:W_{E}\longrightarrow W_{E}^{ab}\simeq E^{\times}\backslash{\mathbb{A}}_{E}^{\times}\longrightarrow{\mathbb{C}}^{\times}$
of the Weil group $W_{E}$ via class field theory. By induction, this gives rise to a 3-dimensional representation
$\rho_{\chi}:={\rm Ind}_{W_{E}}^{W_{k}}\chi\quad\text{of $W_{k}$.}$
We say that $\pi$ is obtained from $\chi$ by automorphic induction if its Hecke-Stakae family $c(\pi)$ is equal to the family of Frobenius classes of $\rho_{\chi}$, and express this as $\pi={\rm AI}_{E/k}(\chi)$.
If $E/k$ is an arbitrary Galois extension with cyclic Galois group of prime order, then the existence of ${\rm AI}_{E/k}(\chi)$ has been shown by Arthur-Clozel [AC] using the twisted trace formula. When $E/k$ is non-Galois cubic, the existence of ${\rm AI}_{E/k}(\chi)$ has been shown by Jacquet-Piatetski-Shapiro-Shalika [JPSS, §14] using the converse theorem. There is also an alternative approach by Kazhdan via exceptional theta correspondence [K].
When $\pi={\rm AI}_{E/k}(\chi)$, we shall see that the adjoint lifting of $\pi$ to ${\rm GL}_{8}$ more directly, without recourse to the twisted trace formula. Indeed, from the viewpoint of Galois representations, one has:
$\rho_{\chi}\otimes\rho_{\chi}^{\vee}={\rm Ind}_{W_{E}}^{W_{k}}\chi\otimes{\rm Ind}_{W_{E}}^{W_{k}}\chi^{-1}={\rm Ind}_{W_{E}}^{W_{k}}\left(\left({\rm Ind}_{W_{E}}^{W_{k}}\chi\right)|_{W_{E}}\otimes\chi^{-1}\right).$
We shall now consider the Galois and non-Galois case separately.
8.3. Galois case
When $E/k$ is Galois, let $\sigma$ be a generator of ${\rm Gal}(E/k)\simeq{\mathbb{Z}}/3{\mathbb{Z}}$ and let $\{\omega_{E/k},\omega_{E/k}^{-1}\}$ be the pair of cubic Hecke characters of $\mathbb{A}$$^{\times}$ associated to $E/k$ by class field theory. Then
$\left({\rm Ind}_{W_{E}}^{W_{k}}\chi\right)|_{W_{E}}\simeq\chi+\chi^{\sigma}+\chi^{\sigma^{2}}$
so that
$\rho_{\chi}\otimes\rho_{\chi}^{\vee}\simeq{\rm Ind}_{W_{E}}^{W_{k}}1\oplus{\rm Ind}_{W_{E}}^{W_{k}}(\chi^{\sigma}/\chi)\oplus{\rm Ind}_{W_{E}}^{W_{k}}(\chi^{\sigma^{2}}/\chi).$
Hence,
${\rm Ad}(\rho_{\chi})\simeq\omega_{E/k}\oplus\omega_{E/k}^{-1}\oplus{\rm Ind}_{W_{E}}^{W_{k}}(\chi^{\sigma}/\chi)\oplus{\rm Ind}_{W_{E}}^{W_{k}}(\chi^{\sigma^{2}}/\chi).$
Now observe that each of the summands above has an automorphic analog and hence we may construct the (weak) adjoint lifting of $\pi$ as an isobaric sum:
(8.1) ${\rm Ad}(\pi)=\omega_{E/k}\boxplus\omega_{E/k}^{-1}\boxplus{\rm AI}_{E/k}(\chi^{\sigma}/\chi)\boxplus{\rm AI}_{E/k}(\chi^{\sigma^{2}}/\chi).$
Thus, we see that in this case, the adjoint lifting of $\pi$ is non-cuspidal, as it contains (at least) two ${\rm GL}_{1}$-summands. Indeed, we have the following result:
Theorem 8.2.
Let $\pi$ be a cuspidal representation of ${\rm GL}_{3}$. Then the following are equivalent:
  • (i)
    $\pi$ is an automorphic induction from some Galois cubic extension;
  • (ii)
    ${\rm Ad}(\pi)$ exists as an automorphic representation of ${\rm GL}_{8}$ and contains a ${\rm GL}_{1}$-summand in its isobaric sum decomposition.
  • (iii)
    $\pi$ is invariant under twisting by some nontrivial Hecke character (necessarily cubic).
More precisely, when the above conditions hold, so that $\pi={\rm AI}_{E/k}(\chi)$ for some Galois cubic extension $E/k$ with associated pair of cubic characters $\{\omega_{E/k},\omega_{E/k}^{-1}\}$, then $\pi$ is invariant under twisting by $\omega_{E/k}^{\pm 1}$ and ${\rm Ad}(\pi)$ contains $\omega_{E/k}^{\pm 1}$ as ${\rm GL}_{1}$-summands. Further, we have the following two possibilities for the isobaric sum decomposition of ${\rm Ad}(\pi)$.
  • (a)
    If $\chi^{\sigma}\cdot\chi^{\sigma^{2}}\neq\chi^{2}$, then the characters $\chi^{\sigma}/\chi$ and $\chi^{\sigma^{2}}/\chi$ are non-invariant under ${\rm Gal}(E/k)$. In this case,
    ${\rm Ad}(\pi)=\omega_{E/k}\boxplus\omega_{E/k}^{-1}\boxplus{\rm AI}_{E/k}(\chi^{\sigma}/\chi)\boxplus{\rm AI}_{E/k}(\chi^{\sigma^{2}}/\chi)$
    with the last two summands cuspidal representations of ${\rm GL}_{3}$ (which are dual of each other). The two ${\rm GL}_{3}$-summands are isomorphic to each other if and only if $(\chi^{2})^{\sigma}=\chi^{2}$. In particular, $\pi$ is an automorphic induction from precisely one Galois cubic field (namely $E/k$).
  • (b)
    If $\chi^{\sigma}\cdot\chi^{\sigma^{2}}=\chi^{2}$, then the characters $\chi^{\sigma}/\chi$ and $\chi^{\sigma^{2}}/\chi$ are invariant under ${\rm Gal}(E/k)$. If $\chi^{\sigma}/\chi=\mu\circ N_{E/k}$ for a Hecke character $\mu$ over $k$, then
    ${\rm Ad}(\pi)=\omega_{E/k}\boxplus\omega_{E/k}^{-1}\boxplus\mu\boxplus\mu\omega_{E/k}\boxplus\mu\omega_{E/k}^{-1}\boxplus\mu^{-1}\boxplus\mu^{-1}\omega_{E/k}\boxplus\mu^{-1}\omega_{E/k}^{-1}.$
    In particular, $\mu$ is a cubic character and ${\rm Ad}(\pi)$ is an isobaric sum of $8$ distinct cubic Hecke characters. Hence, $\pi$ can be expressed as an automorphic induction from four distinct Galois cubic field extensions of $k$.
In particular, the self-dual automorphic representation ${\rm Ad}(\pi)$ is the transfer of an automorphic but non-cuspidal representation of ${\rm SO}_{8}$.
Remark:
(i) If $\pi$ can be obtained from two Galois cubic field extensions $E_{1}$ and $E_{2}$, then $\pi$ is invariant under twisting by $\omega_{E_{1}/k}$ and $\omega_{E_{2}/k}$. Hence $\pi$ is also invariant under twisting by $\omega_{E_{1}/k}\cdot\omega_{E_{2}/k}$ and $\omega_{E_{1}/k}\cdot\omega_{E_{2}/k}^{-1}$. The latter two cubic characters correspond to two other Galois cubic extensions $E_{3}$ and $E_{4}$. The four cubic extensions $E_{i}$ are precisely the four cubic subfields contained in the bi-cyclic-cubic (${\mathbb{Z}}/3{\mathbb{Z}}\times{\mathbb{Z}}/3{\mathbb{Z}}$) extension which is the composition of $E_{1}$ and $E_{2}$. Hence, if $\pi$ is an automorphic induction from more than one Galois cubic extension, it is automatically an automorphic induction from four Galois cubic extensions. Part (b) of the theorem implies that $\pi$ cannot be an automorphic induction from more than four Galois cubic extensions.
(ii) The condition on $\chi$ in (b) of the Theorem is the same condition in [L1, §6.1, Prop. 2 (1)], which implies that the global multiplicity of the L-packet $\mathcal{L}[\pi]$ of ${\rm SL}_{3}$ is $2$. Under the condition in (a), the global multiplicity of $\mathcal{L}[\pi]$ is $1$.
Proof.
Let us first show the equivalence of (i)-(iii):
  • (i) $\Longrightarrow$ (ii): this has been shown by our discussion before the statement of the theorem.
  • (ii) $\Longrightarrow$ (iii): Suppose ${\rm Ad}(\pi)$ contains $\mu$ as a ${\rm GL}_{1}$-summand. Then the partial L-function $L^{S}(s,{\rm Ad}(\pi)\times\mu^{-1})$ has a pole at $s=1$. But one has an identity of partial L-functions:
    $L^{S}(s,\mu^{-1})\cdot L^{S}(s,{\rm Ad}(\pi)\times\mu^{-1})=L^{S}(s,\pi\times(\pi\otimes\mu)^{\vee}).$
    Hence the partial L-function on the RHS also has a pole at $s=1$. By Jacquet-Shalika, this implies that $\pi\simeq\pi\otimes\mu$, as desired. Further, $\mu$ is necessarily cubic, since $\omega_{\pi}=\omega_{\pi\otimes\mu}=\omega_{\pi}\cdot\mu^{3}$, where $\omega_{\pi}$ denotes the central character of $\pi$.
  • (iii) $\Longrightarrow$ (i): this is due to Arthur-Clozel [AC]
This proves the equivalence of (i), (ii) and (iii). Moreover, when these conditions hold, we see by (8.1) that if $\pi={\rm AI}_{E/k}(\chi)$, then $\omega_{E/k}^{\pm 1}$ are ${\rm GL}_{1}$-summands of ${\rm Ad}(\pi)$. Indeed, (8.1) gives the more precise description in (a) and (b) immediately.
8.4. Non-Galois case
We now consider the non-Galois case. If $E/k$ is a non-Galois cubic extension, with discriminant quadratic field extension $K=K_{E}$, then $L=E\cdot K_{E}$ is the Galois closure of $k$ in $\overline{k}$. Hence, $L/k$ is an $S_{3}$-extension, and the cubic extension $L/K$ is Galois. We let $\sigma$ be a generator of ${\rm Gal}(L/K)\simeq{\mathbb{Z}}/3{\mathbb{Z}}$ and note that $L$ contains two other cubic subfields besides $E$, namely
$E_{\sigma}=\{\sigma(e):e\in E\}\quad\text{and}\quad E_{\sigma^{2}}=\{\sigma^{2}(e):e\in E\}.$
In this case, if $\chi$ is a Hecke character of $E$, then
$\left({\rm Ind}_{W_{E}}^{W_{k}}\chi\right)|_{W_{E}}\simeq\chi\oplus{\rm Ind}_{W_{L}}^{W_{E}}(\chi|_{W_{L}}^{\sigma}).$
Here, note that $\chi|_{W_{L}}^{\sigma}$ is defined by $\chi^{\sigma}(w)=\chi(\sigma^{-1}w\sigma)$ for $w\in W_{L}$. To simplify notation, we shall omit $|_{W_{L}}$ in what follows and regard the restriction to $W_{L}$ as implicit. Hence
$\rho_{\chi}\otimes\rho_{\chi}^{\vee}\simeq{\rm Ind}_{W_{E}}^{W_{k}}1\oplus{\rm Ind}_{W_{L}}^{W_{k}}(\chi^{\sigma}/\chi),$
so that
${\rm Ad}(\rho_{\chi})\simeq\rho_{E}\oplus{\rm Ind}_{W_{L}}^{W_{k}}(\chi^{\sigma}/\chi).$
Noting that $\rho_{E}\simeq{\rm Ind}_{W_{K}}^{W_{k}}\omega_{L/K}$ and
${\rm Ind}_{W_{L}}^{W_{k}}={\rm Ind}_{W_{K}}^{W_{k}}\circ{\rm Ind}_{W_{L}}^{W_{K}}$
is a composite of inductions through cyclic extensions of prime degrees, we see that the automorphic analog of ${\rm Ad}(\rho_{\chi})$ can be constructed.
More precisely, if $\pi={\rm AI}_{E/k}(\chi)$, then
(8.3) ${\rm Ad}(\pi)={\rm AI}_{K/k}(\omega_{L/K})\boxplus{\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$
as an automorphic representation of ${\rm GL}_{8}$, where we have written $\chi_{L}:=\chi\circ N_{L/E}$ for the base change of $\chi$ from $E$ to $L$. Henceforth, we will write $\rho_{E}$ to denote its automorphic analog ${\rm AI}_{K/k}(\omega_{L/K})$ as well.
Now one has the following analog of Theorem 8.2:
Theorem 8.4.
Let $\pi$ be a unitary cuspidal representation of ${\rm GL}_{3}$. The following are equivalent:
  • (i)
    $\pi$ can be realized as an automorphic induction from a non-Galois cubic extension.
  • (ii)
    ${\rm Ad}(\pi)$ exists as an automorphic representation of ${\rm GL}_{8}$ and contains a unitary cuspidal ${\rm GL}_{2}$-summand in its isobaric decomposition.
  • (iii)
    for some unitary cuspidal representation $\tau$ of ${\rm GL}_{2}$, $\pi\boxtimes\tau\supset\pi$ in its isobaric decomposition (where $\pi\boxtimes\tau$ is the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ to ${\rm GL}_{6}$ established by Kim-Shahidi [KS1]).
More precisely, if $\pi={\rm AI}_{E/k}(\chi)$ for a non-Galois cubic extension $E/k$ with Galois closure $L/k$ and discriminant quadratic field $K/k$, then ${\rm Ad}(\pi)$ contains the cuspidal representation $\rho_{E}$ of ${\rm GL}_{2}$ as a summand in its isobaric decomposition. Further, we have the following possibilities for the isobaric decomposition of ${\rm Ad}(\pi)$:
  • (a)
    If
    $\chi_{L}^{\sigma}\cdot\chi_{L}^{\sigma^{2}}\neq\chi_{L}^{2}\quad\text{and}\quad(\chi_{L}^{\sigma})^{2}\neq\chi_{L}^{2},$
    then ${\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ is a self-dual cuspidal representation of ${\rm GL}_{6}$ over $k$ with central character $\omega_{K/k}$, so that
    ${\rm Ad}(\pi)=\rho_{E}\oplus{\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right).$
  • (b)
    If
    $(\chi_{L}^{\sigma})^{2}=\chi_{L}^{2},$
    then ${\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ is a self-dual cuspidal representation of ${\rm GL}_{3}$ over $K$ which is invariant under ${\rm Gal}(K/k)$, so that
    ${\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right)=A\oplus A\cdot\omega_{K/k}$
    for some self-dual cuspidal representation $A$ of ${\rm GL}_{3}$ with trivial central character and
    ${\rm Ad}(\pi)=\rho_{E}\oplus A\oplus A\cdot\omega_{K/k}.$
  • (c)
    If
    $\chi_{L}^{\sigma}\cdot\chi_{L}^{\sigma^{2}}=\chi_{L}^{2},$
    then
    ${\rm AI}_{L/K}\left(\chi_{L}^{\sigma}/\chi_{L}\right)=\omega_{1}\boxplus\omega_{2}\boxplus\omega_{3}$
    is the isobaric sum of three distinct cubic characters, defining three cyclic cubic extensions $E_{1}$, $E_{2}$ and $E_{3}$ of $K$, such that $E$ and the $E_{i}$’s are the four cubic extensions in the bi-cyclic-cubic extension $E\cdot E_{i}$ of $K$, and each $E_{i}/k$ is an $S_{3}$-extension. In this case,
    ${\rm Ad}(\pi)=\rho_{E}\oplus\rho_{E_{1}}\oplus\rho_{E_{2}}\oplus\rho_{E_{3}}$
    and $\pi$ can be obtained by automorphic induction from four distinct non-Galois cubic extensions.
In particular, ${\rm Ad}(\pi)$ is the transfer of a globally generic discrete (indeed cuspidal) automorphic representation of ${\rm SO}_{8}$.
Proof.
We first show the equivalence of (i)-(iii).
  • (i) $\Longrightarrow$ (ii): this follows by discussion before the theorem, in particular (8.3).
  • (ii) $\Longrightarrow$ (iii): if ${\rm Ad}(\pi)$ contains a ${\rm GL}_{2}$ cuspidal summand $\tau$ in its isobaric decomposition, then the partial L-function $L^{S}(s,{\rm Ad}(\pi)\times\tau^{\vee})$ has a pole at $s=1$. Now we have the L-function identity:
    $L^{S}(s,\pi\times\pi^{\vee}\times\tau^{\vee})=L^{S}(s,\tau^{\vee})\cdot L^{S}(s,{\rm Ad}(\pi)\times\tau^{\vee}),$
    where all L-functions above are of Langlands-Shahidi type; indeed, the L-function on the LHS occurs when one considers the maximal parabolic subgroup in $E_{6}$ corresponding to the branch vertex in the Dynkin diagram. From this identity, we deduce that the LHS has a pole at $s=1$. But
    $L^{S}(s,\pi\times(\pi\boxtimes\tau)^{\vee}))=L^{S}(s,\pi\times\pi^{\vee}\times\tau^{\vee}).$
    Hence $L^{S}(s,\pi\times(\pi\boxtimes\tau^{\vee}))$ has a pole at $s=1$, which implies that $\pi\otimes\tau$ contains $\pi$ in its isobaric decomposition.
  • (iii) $\Longrightarrow$ (i): We shall make use of a cuspidality criterion for the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ to ${\rm GL}_{6}$ shown by Ramakrishnan-Wang [RW]. Suppose that $\pi\boxtimes\tau$ contains $\pi$ in its isobaric decomposition. By [RW, Thm. 3.1], this can happen only if $\pi$ is an automorphic induction from a non-Galois cubic extension of $k$.
It follows from (8.3) that when $\pi={\rm AI}_{E/k}(\chi)$ for some non-Galois cubic extension $E/k$, then ${\rm Ad}(\pi)$ contains the cuspidal ${\rm GL}_{2}$-summand $\rho_{E}$. To determine the possible isobaric decomposition of ${\rm Ad}(\pi)$, it remains to investigate the term
[…]
We examine the mutually exclusive cases (a)-(c) in the theorem in turn:
  • (a)
    Under the two conditions in (a), ${\rm AI}_{L/K}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ is a cuspidal representation of ${\rm GL}_{3}$ over $K$, and is not invariant under ${\rm Gal}(K/k)$. Hence ${\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ is a self-dual cuspidal representation of ${\rm GL}_{6}$ over $k$, which is self-dual with central character $\omega_{K/k}$ (since ${\rm Ad}(\pi)$ is self-dual with trivial central character).
  • (b)
    Under the conditions in (b), ${\rm AI}_{L/K}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ is a self-dual cuspidal representation of ${\rm GL}_{3}$ over $K$ with trivial central character, which is invariant under ${\rm Gal}(K/k)$. This implies that ${\rm AI}_{L/k}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ has the shape indicated in (b).
  • (c)
    Under the conditions in (c), ${\rm AI}_{L/K}\left(\chi_{L}^{\sigma}/\chi_{L}\right)$ is the isobaric sum of three automorphic characters $\omega_{i}$, satisfying
    $(\omega_{i})_{L}=\chi_{L}^{\sigma}/\chi_{L}.,$
    so that $\omega_{2}=\omega_{1}\cdot\omega_{L/K}$ and $\omega_{3}=\omega_{1}\cdot\omega_{L/K}^{-1}$ (without loss of generality). Each of the $\mu_{i}$’s is a cubic character; this is because $\mu_{i}$ occurs as a summand in the isobaric decomposition of ${\rm Ad}(\pi_{K})$, where $\pi_{K}$ denotes the base change of $\pi$ to $K$, which is still cuspidal. With $E_{i}/K$ denoting the Galois cubic extensions defined by $\omega_{E_{i}/K}$, it is clear from the above relations between $\omega_{i}$ and $\omega_{L/K}$ that $E$ and the $E_{i}$’s are the four cubic subfields in the bi-cyclic-cubic field $E\cdot E_{i}$ (alternatively, one can appeal to Theorem 8.2, applied to $\pi_{K}$. To see that $E_{i}$ is non0Galis over $k$, we observe that if $c$ is the non-trivial element in ${\rm Gal}(L/E)$ (so that $c|_{K}$ is the non-trivial element in ${\rm Gal}(K/k)$), then
    $(\chi_{L}^{\sigma}/\chi_{L})^{c}=\chi_{L}^{\sigma_{2}}/\chi_{L}=(\chi_{L}^{\sigma}/\chi_{L})^{-1},$
    under the hypothesis of (c). This implies that the conclusion of (c).
Corollary 8.5.
A cuspidal representation of ${\rm GL}_{3}$ cannot be simultaneously obtained as an automoprhic induction from a Galois cubic extension and also an automorphic induction from a non-Galois cubic extension.
9. Self-Dual Case
In this section, we consider the case when the cuspidal representation $\pi$ of ${\rm GL}_{3}$ is the twist of a self-dual cuspidal representation. In this case, there is no loss of generality in assuming that $\pi$ is self dual and hence is the adjoint lift (== Gelbart-Jacquet lift) of a cuspidal representation $\tau$ of ${\rm GL}_{2}$ [GJ]. We shall write $\pi={\rm ad}_{2}(\tau)$.
From a Galois theoretic point of view, if $\tau$ is associated with a 2-dimensional Galois representation $\rho$, then ${\rm ad}_{2}(\tau)$ is associated to the 3-dimensional Galois representation ${\rm ad}_{2}(\rho):={\rm Sym}^{2}(\rho)\otimes\det(\rho)^{-1}$. Then
${\rm Ad}({\rm ad}_{2}(\rho))\simeq{\rm ad}_{2}(\rho)\oplus{\rm ad}_{4}(\rho)$
where
${\rm ad}_{4}(\rho)={\rm Sym}^{4}(\rho)\otimes\det(\rho)^{-2}.$
Now the ${\rm Sym}^{4}$-lifting from ${\rm GL}_{2}$ to ${\rm GL}_{5}$ has been shown by H. Kim [Ki], and hence the automorphic analog of ${\rm ad}_{4}(\rho)$ is known to exist; we shall denote it by ${\rm ad}_{4}(\tau)$. In particular, this allows us to construct the adjoint lifting of $\pi={\rm ad}_{2}(\tau)$ as:
(9.1) ${\rm Ad}(\pi)=\pi\boxplus{\rm ad}_{4}(\tau)={\rm ad}_{2}(\tau)\boxplus{\rm ad}_{4}(\tau).$
Now we have:
Theorem 9.2.
Let $\pi$ be a unitary cuspidal representation of ${\rm GL}_{3}$ and assume that $\pi$ is not an automorphic induction from a cubic field extension of $k$ (Galois or not). Then the following are equivalent:
  • (i)
    $\pi$ is the twist by an automorphic character of a self-dual cuspidal representation;
  • (ii)
    ${\rm Ad}(\pi)$ exists as an automorphic representation of ${\rm GL}_{8}$ and contains a cuspidal ${\rm GL}_{3}$-summand in its isobaric decomposition;
More precisely, when the above conditions hold with $\pi$ self-dual, the isobaric decomposition of ${\rm Ad}(\pi)$ is of type $3+5$ and is given by (9.1). In particular, ${\rm Ad}(\pi)$ is the transfer of a globally generic discrete (indeed cuspidal) automorphic representation of ${\rm SO}_{8}$.
Proof.
Assuming (i), then (ii) follows by (9.1). Now assume (ii) and suppose that ${\rm Ad}(\pi)$ contains a cuspidal ${\rm GL}_{3}$-summand $A$ in its isobaric decomposition. By the hypotheses on $\pi$ and Theorems 8.2 and 8.4, we see that the isobaric decomposition of ${\rm Ad}(\pi)$ is of type $3+5$, so that
${\rm Ad}(\pi)=A\boxplus B$
with $B$ a cuspidal ${\rm GL}_{5}$-summand. Since ${\rm Ad}(\pi)={\rm Ad}(\pi)^{\vee}$ is necessarily self-dual with trivial central character, we deduce that $A$ and $B$ are both self-dual with the same central character. As $3$ and $5$ are odd, the self-dual representations $A$ and $B$ must be of orthogonal type.
By Corollary 2.3, we see that
(9.3) $L^{S}(s,{\rm Ad}(\pi)\times{\rm Ad}(\pi))=L^{S}(s,{\rm Ad}(\pi))\cdot L^{S}(s,{\rm Ad}(\pi),\wedge^{3}).$
Here, note that all the partial L-functions in this identity are Langlands-Shahidi L-functions. Indeed, the exterior cube L-function for ${\rm GL}_{8}$ can be obtained as a Langlands-Shahidi L-function by considering the maximal parabolic subgroup of $E_{8}$ with Levi factor of type $A_{7}$. Since ${\rm Ad}(\pi)$ has two cuspidal summands in its isobaric decomposition, the LHS has a pole of order $2$ at $s=1$. Hence $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ has a pole of order $2$ at $s=1$, since $L^{S}(s,{\rm Ad}(\pi))$ is entire and nonvanishing at $s=1$.
On the other hand, as ${\rm Ad}(\pi)=A\oplus B$, $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ is equal to:
(9.4) $L^{S}(s,A,\wedge^{3})\cdot L^{S}(s,A\times B,\wedge^{2}\times{\rm std})\cdot L^{S}(s,A\times B,{\rm std}\times\wedge^{2})\cdot L^{S}(s,B,\wedge^{3}).$
Now we note that:
  • The L-function
    $L^{S}(s,A,\wedge^{3})=L^{S}(s,\omega_{A}),$
    where $\omega_{A}$ is the central character of $A$, has a pole at $s=1$ if and only if $\omega_{A}$ is trivial;
  • the L-function
    $L^{S}(s,A\times B,\wedge^{2}\times{\rm std})=L^{S}(s,(A\cdot\omega_{A})\times B)$
    is holomorphic at $s=1$;
  • the L-function
    $L^{S}(s,A\times B,{\rm std}\times\wedge^{2})$
    has at most a simple pole at $s=1$ by a result of Kim-Shahidi [KS3, Thm. 1(b) and Remark 1].
  • the L-function
    $L^{S}(s,B,\wedge^{3})=L^{S}(s,B,\wedge^{2}\times\omega_{B})$
    is a twisted exterior square L-function and does not have a pole at $s=1$ (since $5$ is odd).
Taken together, we see that the product of L-functions in (9.4) has at most a double pole at $s=1$ and this double pole is achieved ionly if $\omega_{A}=1$. Since we know that $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ has a double pole at $s=1$, we deduce that $\omega_{A}=\omega_{B}$ is trivial. This shows that
$A={\rm ad}_{2}(\tau)$
is the Gelbert-Jacquet lift of a cuspidal representation $\tau$ of ${\rm GL}_{2}$. Moreover, $\tau$ is non-dihedral since $A$ is cuspidal.
It remains to show that $\pi$ is a twist of $A={\rm ad}_{2}(\tau)$ by an automorphic character. For this, we note that the partial L-function $L^{S}(s,{\rm Ad}(\pi)\times A)$ has a pole at $s=1$. Now consider the L-function
$L^{S}(s,(\pi\boxtimes\tau)\times(\pi\boxtimes\tau)^{\vee})$
where $\pi\boxtimes\tau$ is an automorphic representation on ${\rm GL}_{6}$ obtained via the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ shown by Kim-Shahidi [KS1]. This L-function can be expressed as:
$L^{S}(s,{\rm Ad}(\pi)\times{\rm ad}_{2}(\tau))\cdot L^{S}(s,{\rm Ad}(\pi))\cdot L^{S}(s,{\rm ad}_{2}(\tau)\cdot\zeta^{S}(s)$
$=$ $L^{S}(s,{\rm Ad}(\pi)\times A)\cdot L^{S}(s,{\rm Ad}(\pi))\cdot L^{S}(s,A)\cdot\zeta^{S}(s).$
We thus deduce that $L^{S}(s,(\pi\boxtimes\tau)\times(\pi\boxtimes\tau)^{\vee})$ has a double pole at $s=1$; in particular, $\pi\boxtimes\tau$ is not cuspidal on ${\rm GL}_{6}$. Since $\tau$ is not dihedral, it follows by the cuspidal criterion of Ramakrishnan-Wang [RW, Thm. 3.1(a)] that $\pi$ is a twist of $A={\rm ad}_{2}(\tau)$ by an automorphic character, as desired. ∎
Corollary 9.5.
Let $\pi$ be a self-dual cuspidal automorphic representation of ${\rm GL}_{3}$ which is not an automorphic induction from a cubic field extension. If $\pi^{\prime}$ is a cuspidal automorphic representation such that $\pi^{\prime}{\thicksim}_{w}\pi$, then $\pi^{\prime}{\thicksim}_{s}\pi$. Hence, the global multiplicity of the cuspidal L-packet $\mathcal{L}[\pi]$ is $1$.
Proof.
By (9.1) and Theorem 9.2 applied to $\pi$, ${\rm Ad}(\pi)$ is automorphic on ${\rm GL}_{8}$ and has isobaric decomposition $\pi\boxplus\tau$ for some cuspidal ${\rm GL}_{5}$-summand $\tau$. By hypothesis, $\pi^{\prime}$ is nearly equivalent to $\pi\otimes\chi$ for some abstract character $\chi$ of $\mathbb{A}$$^{\times}$. Hence, ${\rm Ad}(\pi^{\prime})$ exists as an automorphic representation on ${\rm GL}_{8}$, since one could simply take
${\rm Ad}(\pi^{\prime}):={\rm Ad}(\pi)=\pi\boxplus\tau.$
Hence, by Theorems 8.2 and 8.4, we see that $\pi^{\prime}$ is also not an automorphic induction from a cubic field extension. Moreover, by Theorem 9.2 applied to $\pi^{\prime}$, we deduce that
$\pi^{\prime}\simeq\pi_{0}\otimes\mu$
for some self-dual cuspidal representation $\pi_{0}$ and an automorphic character $\mu$. But then we would have the isobaric decomposition
${\rm Ad}(\pi^{\prime})=\pi_{0}\boxplus\tau_{0},$
for some ${\rm GL}_{5}$-summand $\tau_{0}$. Comparing the two isobaric decompositions of ${\rm Ad}(\pi^{\prime})$ obtained above, we conclude that $\pi=\pi_{0}$. In particular,
$\pi^{\prime}\simeq\pi\otimes\mu\quad\text{ so that $\pi^{\prime}{\thicksim}_{s}\pi$.}$
Remark: Suppose $\pi={\rm ad}_{2}(\tau)$ is a self-dual cuspidal automorphic representation of ${\rm GL}_{3}$, so that $\tau$ is not dihedral. Then as noted in [KS2, Thm. 3.3.7], one has:
  • $\pi$ is not an automorphic induction from a cyclic cubic extension if and only if $\tau$ is not tetrahedral, in which case ${\rm Sym}^{3}(\tau)$ is cuspidal;
  • $\pi$ is not an automorphic induction from a cubic extension (cyclic or noncyclic) if and only if $\tau$ is neither tetrahedral nor octahedral, in which case both ${\rm Sym}^{3}(\tau)$ and ${\rm ad}_{4}(\tau)$ are cuspidal.
10. Cuspidality Criterion
In this final section, we consider the case when the cuspidal representation $\pi$ of ${\rm GL}_{3}$ is not of the type considered in the previous two sections, i.e. $\pi$ is neither an automorphic induction from a cubic field extension, nor a twist of a self-dual representation. We shall make the following hypothesis:
Hypothesis: Assume that the adjoint lifting ${\rm Ad}(\pi)$ has been constructed as an automorphic representation of ${\rm GL}_{8}$ via the sequence of functorial lifting in (1.1).
By Theorem 7.2, the hypothesis is satisfied when $\pi$ has a discrete series local component. Now one has:
Theorem 10.1.
Let $\pi$ be a unitary cuspidal automorphic representation of ${\rm GL}_{3}$ which is not an automorphic induction from a cubic field extension or a twist of a self-dual representation. Assume that the above hypothesis holds for $\pi$. Then ${\rm Ad}(\pi)$ is a self-dual automorphic representation of orthogonal type with trivial central character. Moreover, one of the following holds:
  • (i)
    ${\rm Ad}(\pi)$ is cuspidal and $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ has a simple pole at $s=1$.
  • (ii)
    ${\rm Ad}(\pi)=A\boxplus B$ with $A$ and $B$ self-dual cuspidal representations of orthogonal type with the same central character which is a nontrivial quadratic character. If $K/k$ is the quadratic field extension determined by the central character of $A$, then $\pi_{K}$ is dihedral with respect to four distinct non-Galois cubic extensions of $K$. Moreover, $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ has a double pole at $s=1$.
Proof.
It is clear that if ${\rm Ad}(\pi)$ exists as an automorphic representation on ${\rm GL}_{8}$, then it is self dual with trivial central character, regardless of whether the adjoint lifting is constructed via the sequence of liftings depicted in (1.1). If it is constructed via (1.1), then it follows that $\pi$ is of orthogonal type, since it is lifted from the split ${\rm SO}_{8}$. This shows the first assertion of the Theorem.
In view of the hypotheses on $\pi$ in the Theorem, as well as Theorems 8.2, 8.4 and 9.2, we already know that ${\rm Ad}(\pi)$ cannot contain any cuspidal ${\rm GL}_{1}$-, ${\rm GL}_{2}$- or ${\rm GL}_{3}$-summands. Hence, the only possibilities for the isobaric decomposition of ${\rm Ad}(\pi)$ are of type $4+4$ or $8$. Assume thus that ${\rm Ad}(\pi)=A\boxplus B$ is an isobaric sum of two cuspidal representations of ${\rm GL}_{4}$. We need to show that $A$ and $B$ are distinct self-dual of orthogonal type with nontrivial central character.
We will again make use of the identity of L-functions:
(10.2) $L^{S}(s,{\rm Ad}(\pi)\times{\rm Ad}(\pi))=L^{S}(s,{\rm Ad}(\pi))\cdot L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$
given in (9.3). Since ${\rm Ad}(\pi)$ is self-dual, we see that the order of pole at $s=1$ of the LHS is:
[…]
On the other hand, if ${\rm Ad}(\pi)=A\boxplus B$, then $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ is equal to the following product:
$L^{S}(s,A^{\vee}\times\omega_{A})\cdot L^{S}(,B^{\vee}\times\omega_{B})\cdot L^{S}(s,A\times B,\wedge^{2}\times{\rm std})\cdot L^{S}(s,B\times A,\wedge^{2}\times{\rm std}),$
where $\omega_{A}$ and $\omega_{B}$ denote the central characters of $A$ and $B$ respectively. The exterior square functorial lifting from ${\rm GL}_{4}$ to ${\rm GL}_{6}$ has been shown by Kim [Ki]. Writing $\wedge^{2}A$ to denote the automorphic representation of ${\rm GL}_{6}$ which is the exterior square functorial lift of $A$, the above product can be written as:
$L^{S}(s,A^{\vee}\times\omega_{A})\cdot L^{S}(,B^{\vee}\times\omega_{B})\cdot L^{S}(s,\wedge^{2}A\times B)\cdot L^{S}(s,\wedge^{2}B\times A).$
Now the Rankin-Selberg L-functions $L^{S}(s,\wedge^{2}A\times B)$ and $L^{S}(s,\wedge^{2}B\times A)$ have at most simple poles at $s=1$, since the isobaric decomposition of $\wedge^{2}A$ and $\wedge^{2}B$ can contain at most one ${\rm GL}_{4}$-cuspidal summand. Hence, the RHS of (10.2) has at most a double pole at $s=1$. Comparing with (10.3), we conclude that $A$ and $B$ are distinct, and
$\wedge^{2}A\supset B^{\vee}\quad\text{and}\quad\wedge^{2}B\supset A^{\vee}$
in their isobaric decompositions. In particular, $\wedge^{2}A$ and $\wedge^{2}B$ are not cuspidal and contain a cuspidal ${\rm GL}_{4}$-summand.
We shall show:
Lemma 10.4.
Assume that $A$ and $B$ are distinct cuspidal representations of ${\rm GL}_{4}$ such that $\wedge^{2}A\supset B^{\vee}$ and $\wedge^{2}B\supset A^{\vee}$. Then at least one of $A$ or $B$ is invariant under twisting by a nontrivial quadratic character.
Proof.
We will make use of the cuspidality criterion for the exterior square lifting of ${\rm GL}_{4}$ established by Asgari and Raghuram [AR]. By [AR, Prop. 3.1, Prop. 3.2, Prop. 3.3 and Prop. 3.5], we see that for $\wedge^{2}A$ to contain a cuspidal ${\rm GL}_{4}$-summand in its isobaric decompostion, one of the following must hold:
  • (a)
    $A\simeq{\rm AI}_{K/k}(\tau)$ is an automorphic induction of a cuspidal representation $\tau$ of ${\rm GL}_{2}$ over a quadratic field extension $K$ of $k$;
  • (b)
    $A\simeq{\rm As}_{K/k}^{+}(\tau)$ is an Asai lift of a cuspidal representation $\tau$ of ${\rm GL}_{2}$ over a quadratic field extension $K$ of $k$;
  • (c)
    AA is obtained as a functorial lift for a globally generic cuspidal representation $\sigma$ of ${\rm GSp}_{4}$.
If (a) holds, then $A$ is invariant under twisting by the nontrivial quadratic character $\omega_{K/k}$ and we are done. We examine the remaining two cases in turn:
  • (b)
    Under (b), it follows by [AR, Proof of Prop. 3.2] that
    $\wedge^{2}A={\rm AI}_{K/k}({\rm ad}_{2}(\tau))\cdot\omega_{\tau}|_{{\mathbb{A}}_{k}^{\times}}.$
    For this to contain a cuspidal ${\rm GL}_{4}$-summand, it is necessary that ${\rm ad}_{2}(\tau)$ has isobaric decomposition of type $1+2$, say ${\rm ad}_{2}(\tau)=\nu\boxplus\sigma$, in which case the cuspidal ${\rm GL}_{4}$-summand is
    $B^{\vee}={\rm AI}_{K/k}(\sigma)\cdot\omega_{\tau}|_{{\mathbb{A}}_{k}^{\times}}.$
    This shows that $B$ is invariant under twisting by the nontrivial quadratic character $\omega_{K/k}$.
  • (c)
    Under (c), $A$ is the functorial lift of a globally generic cuspidal representation $\sigma$ of ${\rm GSp}_{4}$. In fact, this functorial lift can be constructed as a similitude theta lift from ${\rm GSp}_{4}$ to ${\rm GSO}_{6}$ (which is a quotient of ${\rm GL}_{4}\times{\rm GL}_{1}$ by a central $\mathbb{G}_{m}$). By [AR, Prop. 3.3],
    $\wedge^{2}A={\rm std}_{*}(\sigma|_{{\rm Sp}_{4}})\cdot\omega_{\sigma}\boxplus\omega_{\sigma},$
    where ${\rm std}_{*}(\sigma)$ is the automorphic representation of ${\rm GL}_{5}$ obtained as a functorial lift of $\sigma$ via the morphism of dual groups
    ${\rm std}:{\rm GSp}_{4}^{\vee}={\rm GSp}_{4}({\mathbb{C}})\longrightarrow{\rm PGSp}_{4}({\mathbb{C}})\simeq{\rm SO}_{5}({\mathbb{C}})\hookrightarrow{\rm GL}_{5}({\mathbb{C}}).$
    The functorial lifting for the first morphism is given simply by the restriction of automorphic forms from ${\rm GSp}_{4}$ to ${\rm Sp}_{4}$, whereas the second was shown by [CKPSS, A].
    For $\wedge^{2}A$ to contain a cuspidal ${\rm GL}_{4}$-summand, it is necessary that ${\rm std}_{*}(\sigma)$ has isobaric decomposition of type $1+4$. Since $B^{\vee}\subset\wedge^{2}A$, we must have ${\rm std}_{*}(\sigma)=\nu\boxplus B^{\vee}\cdot\omega_{\sigma}^{-1}$ for some automorphic quadratic character $\nu$. Now this implies that $L^{S}(s,\sigma|_{{\rm Sp}_{4}},{\rm std}\times\nu)$ has a pole at $s=1$, which implies that $\sigma$ has a nonzero theta lift to a quasi-split group ${\rm GSO}_{4}$ associated to $\nu$. Since $\sigma$ has cuspidal theta lift to the split ${\rm GSO}_{6}$, it does not lift to the split ${\rm GSO}_{4}$ under the theta correspondence. Hence, $\nu$ must be nontrivial and $\sigma$ has nonzero theta lift to a quasi-split but non-split ${\rm GSO}_{4}$ associated to $\nu$. This implies that $A$ is obtained as an automorphic induction from the quadratic field determined by $\nu$ and hence is invariant under twisting by $\nu$.
We have thus completed the proof of the lemma.
With the lemma in hand, we can now continue with the proof of Theorem 10.1. By the lemma, we may assume without loss of generality that $A\otimes\omega_{K/k}\simeq A$ for some quadratic field extension $K/k$, so that $A\simeq{\rm AI}_{K/k}(\tau)$ for some cuspidal representation $\tau$ of ${\rm GL}_{2}$ over $K$. Consider now the base change $\pi_{K}$ of $\pi$ to $K$, noting that $\pi_{K}$ is still cuspidal. Then
${\rm Ad}(\pi_{K})={\rm Ad}(\pi)_{K}=A_{K}\boxplus B_{K}=\tau\boxplus\tau^{c}\boxplus B_{K}$
where $c$ is the nontrivial element in ${\rm Gal}(K/k)$ and the base change $B_{K}$ has isobaric decomposition of type $2+2$ or $4$.
Now we may appeal to Theorem 8.4. This says that:
  • $\tau=\rho_{E}$ for some non-Galois cubic extension $E/K$, whose Galois closure $\tilde{E}$ is an $S_{3}$-extension of $K$ with discriminant quadratic field $L/K$, so that
    $A_{K}=\rho_{E}\oplus\rho_{E^{c}}.$
  • the bi-cyclic-cubic field extension $\tilde{E}\cdot\tilde{E}^{c}$ over $L$ contains two other cubic extensions $\tilde{E}_{1}$ and $\tilde{E}_{2}=\tilde{E}_{1}^{c}$, which are $S_{3}$-extensions over $K$ with discriminant field extension $L/K$, so that
    $B_{K}=\rho_{E_{1}}\boxplus\rho_{E_{2}},$
    where $E_{1}$ and $E_{2}$ are cubic subfields of $\tilde{E}_{1}/K$ and $\tilde{E}_{2}/K$.
In partiular, we have:
$A\simeq{\rm AI}_{K/k}(\rho_{E})\quad\text{and}\quad B\simeq{\rm AI}_{K/k}(\rho_{E_{1}}).$
Since $\rho_{E}$ and $\rho_{E_{1}}$ are self-dual representations with central character $\omega_{L/K}$, we deduce that $A$ and $B$ are self-dual and
$\omega_{A}=\omega_{B}=\omega_{L/K}|_{{\mathbb{A}}^{\times}}.$
On the other hand, with $A$ and $B$ as above, one has [AR, proofs of Prop. 3.2 and Prop. 3.5]
$\wedge^{2}A={\rm As}_{K/k}^{+}(\rho_{E})\cdot\omega_{K/k}\boxplus{\rm AI}_{K/k}(\omega_{L/K})\quad\text{and}\quad\wedge^{2}B={\rm As}^{+}_{K/k}(\rho_{E_{1}})\cdot\omega_{K/k}\boxplus{\rm AI}_{K/k}(\omega_{L/K}),$
so that
$A\simeq{\rm As}^{+}_{K/k}(\rho_{E_{1}})\cdot\omega_{K/k}\quad\text{and}\quad B\simeq{\rm As}^{+}_{K/k}(\rho_{E})\cdot\omega_{K/k}.$
From these alternate expressions of $A$ and $B$, one deduces by [Ra2] that
$\omega_{A}=\omega_{K/k}.$
In particular, since $\omega_{A}$ and $\omega_{B}$ are nontrivial, $A$ and $B$ are necessarily of orthogonal type. Moreover, we also conclude that $\omega_{L/K}|_{{\mathbb{A}}^{\times}}=\omega_{K/k}$, which implies that $L/k$ is a cyclic degree 4 extension. ∎
Here is an interesting corollary, which is an application of our results on the possible isobaric decompositions for ${\rm Ad}(\pi)$ to the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ to ${\rm GL}_{6}$.
Corollary 10.5.
Suppose that
  • $\pi$ is a unitary cuspidal representation of ${\rm GL}_{3}$ which is neither an automorphic induction from a cubic field extension nor a twist of a self-dual cuspidal representation, and ${\rm Ad}(\pi)$ exists on ${\rm GL}_{8}$ (for example when $\pi$ has a discrete series local component);
  • $\tau_{1}$ and $\tau_{2}$ are cuspidal representations of ${\rm GL}_{2}$ with the same central character.
Then
$\pi\boxtimes\tau_{1}\simeq\pi\boxtimes\tau_{2}\Longrightarrow\tau_{1}\simeq\tau_{2}.$
Proof.
If $\pi\boxtimes\tau_{1}\simeq\pi\boxtimes\tau_{2}$, then the partial L-function
$L^{S}(s,(\pi\boxtimes\tau_{1})\times(\pi\boxtimes\tau_{2})^{\vee})=L^{S}(s,\pi\times\pi^{\vee}\times\tau_{1}\times\tau_{2}^{\vee})$
has a pole at $s=1$. But this L-function is equal to
$L^{S}({\rm Ad}(\pi)\times(\tau_{1}\boxtimes\tau_{2}^{\vee}))\cdot L^{S}(s,\tau_{1}\times\tau_{2}^{\vee}).$
Under the hypothesis on $\pi$, it follows by Theorem 10.1 that ${\rm Ad}(\pi)$ is either cuspidal or an isobaric sum of two self-dual cuspidal representations of ${\rm GL}_{4}$ with nontrivial central character. This implies that the L-function $L^{S}({\rm Ad}(\pi)\times(\tau_{1}\boxtimes\tau_{2}^{\vee}))$ is holomorphic at $s=1$, since the self-dual representation $\tau_{1}\boxtimes\tau_{2}$ has trivial central character. Hence, the L-function $L^{S}(s,\tau_{1}\times\tau_{2}^{\vee})$ must have a pole at $s=1$, which implies that $\tau_{1}\simeq\tau_{2}$. ∎
10.1. Summary
The following proposition summarizes an interesting property which should play an important role in the characterization of the image of the adjoint lifting.
Corollary 10.6.
Assume that $\pi$ is a cuspidal automorphic representation of ${\rm GL}_{3}$ which is not invariant under twisting by a nontrivial cubic automorphic character. If $\pi$ is in the context of Theorem 10.1, assume that ${\rm Ad}(\pi)$ is constructed via the sequence of lifting (1.1). Then ${\rm Ad}(\pi)$ is a discrete orthogonal L-parameter. Moreover,
${\rm ord}_{s=1}L^{S}(s,{\rm Ad}(\pi),\wedge^{3})={\rm ord}_{s=1}L^{S}(s,{\rm Ad}(\pi),{\rm Sym}^{2})<0,$
so that the order of poles at $s=1$ of $L^{S}(s,{\rm Ad}(\pi),\wedge^{3})$ (which is an L-function of Langlands-Shahidi type) is equal to the number of summands in the isobaric decomposition of ${\rm Ad}(\pi)$.
One may wonder if the above property of the exterior cube L-function characterizes those discrete orthogonal parameter which arises as the adjoint lifting from ${\rm GL}_{3}$. As the exterior cube L-function occurs as a Langlands-Shahidi L-function in the group $E_{8}$, it is conceivable that one may establish this expectation by an automorphic decent construction from ${\rm GL}_{8}$ to ${\rm GL}_{3}$.
11. Heisenberg-Weil Representations
In Theorems 8.2, 8.4, 9.2 and 10.1, we have described the possible shapes of the isobaric decomposition of ${\rm Ad}(\pi)$. One may ask if all these possibilities are actually attained. While it is not so surprising that the more nondegenerate cases actually happen, it may be less clear for the most degenerate cases in Theorem 8.2(b), Theorem 8.4(c) and Theorem 10.1(ii). In this final section, we will give a general construction of cuspidal representations $\pi$ such that ${\rm Ad}(\pi)$ has the form in these degenerate cases. This construction was first studied by Blasius [B] and Lapid [L2, §5] in their work on multiplicities and rigidity of cuspidal $L$-packets of ${\rm SL}_{n}$.
11.1. Jacobi group
We shall consider the two dimensional symplectic vector space $W$ over the finite field ${\mathbb{Z}}/3{\mathbb{Z}}$. Choosing a Witt basis for $W$, we may identify $W$ with ${\mathbb{Z}}/3{\mathbb{Z}}\times{\mathbb{Z}}/3{\mathbb{Z}}$. Let ${\rm Sp}(W)$ denote the finite symplectic group attached to $W$, which is isomorphic to ${\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})$. Note that
${\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})\simeq Q_{8}\rtimes{\mathbb{Z}}/3{\mathbb{Z}}$
where $Q_{8}$ denotes the finite quaternion group (with $8$ elements). In particular, $|{\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})|=24$ and $Q_{8}$ is its unique $2$-Sylow subgroup. Indeed, $Q_{8}$ is also the commutator subgroup of ${\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})$.
One can also form the Heisenberg group
$H(W)=W\oplus Z,\quad\text{with center $Z={\mathbb{Z}}/3{\mathbb{Z}}$.}$
Via its natural action on $W$, ${\rm Sp}(W)$ acts on $H(W)$ as group automorphisms fixing the center $Z$ elementwise. Hence one has the Jacobi group
$J(W)=H(W)\rtimes{\rm Sp}(W).$
For our purpose, we will consider the following sequence of subgroups of $J(W)$:
(11.1) $H(W)\subset H(W)\cdot Z_{{\rm Sp}(W)}\subset H(W)\cdot C_{4}\subset H(W)\cdot Q_{8}\subset J(W),$
where $Z_{{\rm Sp}(W)}\simeq\mu_{2}$ is the center of ${\rm Sp}(W)$ and $C_{4}$ is the cyclic group of order $4$ generated by the Weyl element
$w_{0}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$
11.2. Heisenberg-Weil representation
By the Stone-von-Neumamn theorem, for a given nontrivial central character $\chi:{\mathbb{Z}}/3{\mathbb{Z}}\rightarrow{\mathbb{C}}^{\times}$, there is a unique (up to isomorphism) irreducible representation $\omega_{\chi}$ of $H(W)$. This irreducible representation is of dimension $3$ and can be constructed as follows. Let $L\subset W$ be any line (i.e. one-dimensional subspace) in $W$. Then $H(L):=L\oplus Z\subset H(W)$ is an abelian subgroup and we may extend $\chi$ from $Z$ to $H(L)$ trivially on $L$. Denoting this extended character by $\chi_{L}$, one has
$\omega_{\chi}={\rm Ind}_{H(L)}^{H(W)}\chi_{L}.$
By the theory of Weil representations, this representation can be extended to to ${\rm Sp}(W)$. Since ${\rm Sp}(W)={\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})$ is not perfect, the extension is of course not unique, as we may twist by cubic characters of ${\rm Sp}(W)$. However, the restriction of these extensions to the commutator subgroup $Q_{8}$ are isomorphic. We will denote this extended representation of $H(W)\rtimes Q_{8}$ by $\omega_{\chi}$ as well, and call it the Heisenberg-Weil representation. The restriction of $\omega_{\chi}$ to each subgroup of $H(W)\rtimes Q_{8}$ is thus irreducible.
11.3. Setting of Theorem 8.2
Let $M/F$ be a Galois extension with Galois group isomorphic to the Heisenberg group $H(W)$ and consider its irreducible 3-dimensional representation $\omega_{\chi}$ constructed above. The fixed field of the center $Z$ of $H(W)$ is thus a bi-cyclic-cubic extension $E_{Z}$ of $F$. For any line $L\subset W$, the fixed field of the abelian normal subgroup $H(L)$ is a Galois cubic extension $E_{L}$ of $F$. Since
$\omega_{\chi}\simeq{\rm Ind}_{H(L)}^{H(W)}\chi_{L},$
we see that $\omega_{\chi}$ gives rise to the cuspidal representation
$\pi={\rm AI}_{E_{L}/F}(\chi_{L}).$
In other words, $\pi$ is an automorphic induction from $E_{L}$. Since there are 4 possible lines $L\subset W$ (as $|\mathbb{P}^{1}(\mathbb{F}_{3})|=4$), we see that $\pi$ is an automorphic induction from the 4 cubic subfields in the bi-cyclic-cubic extension $E_{Z}$ of $F$, so that we are in situation (b) of Theorem 8.2. In particular, ${\rm Ad}(\pi)$ is the sum of the 8 cubic characters $\omega_{E_{L}/F}^{\pm 1}$.
11.4. Setting of Theorem 8.4
Now suppose that $M/F$ is a Galois extension with Galois group $H(W)\rtimes\mu_{2}$. Let $K$ be the fixed field of $H(W)$, so that $K/F$ is a quadratic extension. The action of ${\rm Gal}(K/F)$ on Gal(M/K){Gal}(M/K) is the action of $\mu_{2}\subset{\rm Sp}(W)$ on $H(W)$. In particular, it acts as 1-1 on ${\rm Gal}(E_{Z}/K)=W$ (where $E_{Z}$ is the fixed field of the center $Z$ of $H(W)$). This shows that for each line $L\subset W$, its fixed field $E_{L}$ is a cubic extension of $K$, which is Galois over $F$ with ${\rm Gal}(E_{L}/F)\simeq S_{3}$.
Now the irreducible 3-dimensional representation $\omega_{\chi}$ of $H(W)$ gives rise to a cuspidal representation $\pi_{K}$ of ${\rm GL}_{3}({\mathbb{A}}_{K})$ as described above. Hence, $\pi_{K}$ is an automorphic induction from any of the 4 cubic subfields $E_{L}$ over $K$. Since $\omega_{\chi}$ extends to $H(W)\rtimes\mu_{2}$, the cuspidal representation $\pi_{K}$ is invariant under the action of ${\rm Gal}(K/F)$; in particular, it follows from [AC] that $\pi_{K}$ is the base change of a cuspidal representation $\pi$ of ${\rm GL}_{3}$ over $F$. Moreover, since ${\rm Ad}(\pi_{K})$ is the isobaric sum of the 8 cubic characters $\omega_{E_{L}/K}^{\pm 1}$, we deduce that ${\rm Ad}(\pi)$ is the isobaric sum of 4 cuspidal representations of ${\rm GL}_{2}$. In particular, we are in the situation of Theorem 8.4(c).
11.5. Setting of Theorem 10.1
Finally, suppose that $M/F$ is a Galois extension with Galois group $H(W)\rtimes C_{4}$ and set
[…]
Then $L/F$ is a cyclic degree $4$ extension with Galois group $C_{4}$. By the previous subsection, the irreducible representation $\omega_{\chi}$ of $H(W)\rtimes\mu_{2}$ gives rise to a cuspidal representation $\pi_{K}$ of ${\rm GL}_{3}({\mathbb{A}}_{K})$ such that ${\rm Ad}(\pi_{K})$ is the isobaric sum of 4 cuspidal representations of ${\rm GL}_{2}({\mathbb{A}}_{K})$, corresponding to the 4 lines in $W$.
Since $\omega_{\chi}$ extends to $H(W)\rtimes C_{4}$, we see that $\pi_{K}$ is invariant under ${\rm Gal}(K/F)$, and hence is the base change of a cuspidal representation of ${\rm GL}_{3}$ over $F$ by [AC]. Moreover, the action of ${\rm Gal}(L/F)=C_{4}$ on ${\rm Gal}(E_{Z}/L)=W$ (where $E_{Z}$ is the fixed field of the center $Z\subset H(W)$) is generated by the action of the Weyl element
$w_{0}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right).$
As the element $w_{0}^{2}$ acts as 1-1 on $W$, we see that the action of $C_{4}$ has two orbits on the set of lines in $W$. This shows that ${\rm Gal}(K/F)$ has two orbits on the 4 isobaric summands in ${\rm Ad}(\pi_{K})$, and hence ${\rm Ad}(\pi)$ is the isobaric sum of two cuspidal representations of ${\rm GL}_{4}$ over $F$. In particular, we are in the situation of Theorem 10.1(ii).
Not surprisingly, if we consider a Galois extension with Galois group $H(W)\rtimes Q_{8}$, a similar argument as above will lead to a cuspidal $\pi$ such that ${\rm Ad}(\pi)$ is cuspidal on ${\rm GL}_{8}$ (i.e. the situation of Theorem 10.1(i)) but the base change of ${\rm Ad}(\pi)$ to an appropriate quadratic extension decomposes as an isobaric sum of two cuspidal ${\rm GL}_{4}$-summands (i.e. the situation of Theorem 10.1(ii)).
We note that in closing that, by a result of Lapid [L2, §5, Prop. 2], the existence of the adjoint lifting as shown in Theorem 7.2 implies that the Galois representation $\omega_{\chi}$ of $H(W)\rtimes\mu_{2}$, $H(W)\rtimes C_{4}$ or $H(W)\rtimes Q_{8}$ corresponds to the cuspidal representation $\pi$ in the respective cases under the purported global Langlands correspondence.